### v1

# Programming with Pixels: Computer-Use Meets Software Engineering

Abstract
Recent advancements in software engineering (SWE) agents have largely followed a tool-based paradigm, where agents interact with hand-engineered tool APIs to perform specific tasks. While effective for specialized tasks, these methods fundamentally lack generalization, as they require predefined tools for each task and do not scale across programming languages and domains. We introduce Programming with Pixels (PwP), an agent environment that unifies software development tasks by enabling computer-use agents—agents that operate directly within an IDE through visual perception, typing, and clicking, rather than relying on predefined tool APIs. To systematically evaluate these agents, we propose PwP-Bench, a benchmark that unifies existing SWE benchmarks spanning tasks across multiple programming languages, modalities, and domains under a task-agnostic state and action space. Our experiments demonstrate that general-purpose computer-use agents can approach or even surpass specialized tool-based agents on a variety of SWE tasks without the need for hand-engineered tools. However, our analysis shows that current models suffer from limited visual grounding and fail to exploit many IDE tools that could simplify their tasks. When agents can directly access IDE tools, without visual interaction, they show significant performance improvements, highlighting the untapped potential of leveraging built-in IDE capabilities. Our results establish PwP as a scalable testbed for building and evaluating the next wave of software engineering agents. [^1] Carnegie Mellon University. We release code and data at programmingwithpixels.com.
Machine Learning, ICML
1 Introduction
Figure 1: Comparison between the traditional tool-based paradigm (left) and our proposed Programming with Pixels (PwP) framework (right) for software engineering (SWE) agents. Instead of relying on specialized hand-engineered tools, PwP enables agents to interact directly with an IDE through basic computer interactions and screen observations. The framework naturally allows agents to take advantage of existing IDE capabilities across multiple languages and tools. Furthermore, PwP-Bench is a comprehensive benchmark to evaluate agent performance across different SWE domains.
Human software developers possess a remarkable ability to work across a wide range of programming tasks, seamlessly adapting to new languages, tools, and problem domains. Realizing a single, general-purpose agent with similar versatility is the overarching goal of many recent efforts in code generation and software engineering automation (Jiang et al., 2024; Jin et al., 2024; Wang et al., 2024b). However, most software engineering agents (SWE agents) still rely on a tool-based paradigm, where an agent takes actions using hand-engineered functions (e.g., search repository, run Python code) exposed through a text API (Yang et al., 2024a, b; Wang et al., 2024a, b). This fundamentally limits generalization, since tool-based agents can only perform tasks using the predefined actions. For example, an agent designed to manage GitHub pull requests lacks debugging abilities unless it is programmed into the agent’s API. Furthermore, the tool-based paradigm lacks scalability, as hand-engineering complex tools requires significant human effort and may not be bug-free. As a result, it remains unclear whether the tool-based paradigm scales to the diversity of software engineering tasks, which spans multiple languages, modalities, and task types.
Our motivating hypothesis is that achieving general-purpose SWE agents requires a shift to computer-use agents (Anthropic, 2024) that interact with computers as humans do: by observing the screen, typing, and clicking. To this end, we recast agentic software engineering as interacting directly with an integrated development environment (IDE) by observing its visual state and using basic actions such as clicking and typing. This allows the agent to perform any task possible in an IDE and leverage all of the IDE’s tools—from debuggers to web browsers—without requiring specialized APIs. However, despite promising results in web navigation (Anthropic, 2024) and open-ended computer tasks (Xie et al., 2024), the ability of computer-use agents to perform software engineering remains underexplored and we lack a dedicated environment for software engineering.
To close this gap, we introduce Programming with Pixels (PwP), the first software engineering agent environment aimed at general-purpose computer-use agents. The PwP environment is a VSCode-based IDE where agents perceive the screen and use primitive actions such as typing, pointing, and clicking. PwP fulfills two key properties. First, the environment is expressive, allowing agents to complete any software engineering task achievable in an IDE, without language- or domain-specific modifications. Second, agents naturally interact with any tools available in the IDE–including debuggers, linters, and code suggestions while handling diverse data types such as images, videos, and PDFs–through basic actions such as clicking and typing. This notion of tool use fundamentally differs from hand-engineered tool APIs, offering scalability and reducing the effort needed to hand-engineer tools for AI agents. Namely, PwP lets agents take advantage of the rich tools already accessible to humans in an IDE, rather than reinventing the wheel. Finally, computer-use agents reduce the need for complex tool pipelines (e.g., tool-specific prompts), opening up a simplified approach to general-purpose SWE agents.
To further evaluate agents developed for computer-use we construct PwP-Bench, a unified benchmark of 15 tasks spanning a variety of software engineering activities, including code generation, pull request resolution, UI development, and DevOps workflows. We show for the first time that general-purpose computer-use agents achieve non-trivial performance on a wide variety of SWE tasks, often approaching or exceeding state-of-the-art tool-based agents.
However, our analysis reveals substantial opportunities for future work. First, even state-of-the-art computer-use agents suffer from visual grounding issues. Second, we show that current agents lack the ability to use many of the tools available in the IDE, including ones that could make their tasks trivial. This suggests that training agents to explore and use the tools available in the IDE is a fruitful future direction. Finally, we find that only one model, Claude (Anthropic, 2024), performs well, highlighting the need for further research into training and improving computer-use agents.
In summary, our contributions are as follows. First, we introduce Programming with Pixels (PwP), the first software engineering-focused environment for evaluating computer-use agents. Second, we introduce PwP-Bench, a benchmark spanning 15 diverse SWE domains, allowing for systematic comparison of software engineering agents. Third, we demonstrate, for the first time, that computer-use agents can perform a wide variety of software engineering tasks without additional hand-engineering of their action or observation spaces. Fourth, we analyze the limitations of current computer-use agents, identifying the need for models that better leverage IDE tooling and highlighting agent training as a key future direction. Finally, both existing agents and benchmarks can be easily incorporated into our unified environment, positioning PwP to serve as a common platform for developing future SWE agents. Overall, PwP challenges the prevailing tool-based paradigm for SWE agents and provides a platform for developing more general agents that interact directly with IDEs.
2 Related Work
2.1 Task-specific SWE benchmarks
Early neural code generation approaches were typically evaluated on fixed input-output pairs—for example, generating code from docstrings (Chen et al., 2021) or from general textual descriptions (Austin et al., 2021). Subsequent benchmarks extended these evaluations to interactive settings, such as resolving GitHub pull requests or writing unit tests for real-world code repositories (Jimenez et al., 2023; Zan et al., 2024; Mündler et al., 2025). More recently, efforts have broadened the scope of code generation to include multimodal tasks, where vision models must interpret images to generate correct code or edits (Si et al., 2024; Shi et al., 2024; Jing et al., 2024; Yang et al., 2024b). However, each of these benchmarks is confined to specific languages, modalities, or task types. In contrast, our proposed PwP-Bench unifies these diverse evaluations into a single framework, encompassing multimodal and multilingual challenges that require interaction with a broad suite of IDE tools. Using this unified approach we reproduce the performance of established benchmarks and encourage the development of general-purpose agents capable of handling a variety of new software engineering tasks. We further compare our work with previous efforts in Tables 1 and 2.
2.2 Software Engineering (SWE) Agents
Recent work has explored “code agents” that move beyond single-step neural code generation toward interactive methods, where intermediate feedback from tools informs subsequent actions. However, many of these approaches specialize in particular tools or programming languages (Jin et al., 2024; Yang et al., 2024b), limiting their broader applicability. For example, Agentless (Xia et al., 2024) relies on a tool that parses files into Python-specific class and function structures. This fails to perform well in other languages or settings (Yang et al., 2024b) without manual modifications. Similarly, the SWE-agent requires modifications to adapt to different tasks (Abramovich et al., 2024; Yang et al., 2024b). In contrast, agents designed for PwP are inherently task and language-agnostic due to the expressive action and observation spaces mandated by our environment. Moreover, the diverse tasks in PwP-Bench require agents to generalize across a wide range of SWE challenges rather than excel in one narrowly defined area such as resolving pull requests.
Many existing agents also depend on hand-engineered tools that require human effort to implement and are susceptible to bugs. For instance, Agentless (Xia et al., 2024) leverages tools for parsing files into Python-specific structures; CodeAct relies on an IPython kernel (Wang et al., 2024a); SWE-Agent uses dedicated search and file editing tools (Yang et al., 2024a); AutoCodeRover requires a linter (Zhang et al., 2024); SWE-Agent EnIGMA develops specialized tools for CTF-style competitions (Abramovich et al., 2024); and SWE-Bench-MM (Yang et al., 2024b) implements a browser view. In PwP, these tools are inherently available within the IDE (as detailed in Table 7), and the agent’s task is to effectively use them rather than being explicitly guided on which tool to use for each specific task.
Finally, current approaches often blur the line between the agent and the environment, as each agent is designed with its own specified action and observation spaces within a self-created environment. Programming with Pixels addresses this issue by unifying existing environments into a single, general-purpose platform on which agents operate. This clear separation of environment design from agent design standardizes evaluation and also allows any existing agent to be modeled within our framework, making it an important testbed for both current and future SWE agents.
2.3 Visual Agents and Computer-Use Agents
A family of recent multimodal agent benchmarks require agents to operate user interfaces using a predefined, limited set of actions (e.g., new_tab, go_back, click [element id](Koh et al., 2024a; Deng et al., 2023; Zheng et al., 2024) . These visual agents typically rely on additional prompting—such as set-of-marks techniques that supply an HTML accessibility tree containing textual and positional information—to overcome their inherent poor visual grounding capabilities (Yang et al., 2023b). Despite such aids, these agents often fail when faced with the complex and dense IDE interfaces found in our environment.
A separate family of computer-use agents (Anthropic, 2024; OpenAI, 2025; Gou et al., 2024) are trained to operate with an expressive action and observation space using primitive operations like clicks and keystrokes, without the need for external accessibility elements. However, there is no SWE-specific environment for evaluating and further training these agents. PwP fills this gap by providing a unified, expressive IDE platform that challenges computer-use agents with realistic and diverse SWE tasks.
2.4 Expressive Agent Environments
Prior work on expressive agent environments has predominantly targeted the web domain (Koh et al., 2024a; Deng et al., 2023), entire operating systems (Xie et al., 2024; Bonatti et al., 2024; Rawles et al., 2023), or other general scenarios (Xu et al., 2024). Some of these environments, such as OSWorld (Xie et al., 2024), feature general action and observation spaces similar to ours. However, although these benchmarks are capable of expressing a wide range of tasks, they do not focus on the unique challenges inherent to software engineering within an IDE. For example, while OSWorld offers a broad set of tasks, it is not specifically designed for SWE, resulting in increased computational overhead. Software engineering is a diverse and important domain that merits its own dedicated environment.
Additionally, we design PwP so that existing tool-based software engineering agents can be readily incorporated into our framework. Specifically, we modify the source code of the IDE to open up API calls that let us test current tool-based agents. Furthermore, PwP-Bench is tailored specifically for multimodal SWE tasks within an IDE, encompassing activities such as pull-request handling, debugging, and image-based code generation across multiple programming languages. We also observe that existing agents built for generic UI control often struggle in the PwP environment, as they must interact with a richer set of tools and achieve precise visual grounding within a complex interface containing a large number of interactive elements. We further distinguish PwP from other environments in Table 1.
3 Programming with Pixels (PwP)
Modern software engineering (SWE) requires using multiple programming languages, tools, and modalities. Furthermore, it relies on a wealth of tools that required tremendous human effort to create—from linters, to visual code debuggers, to project management tools. Motivated by these observations, we create Programming with Pixels, an IDE environment that satisfies two properties: (i) it is expressive, meaning that an agent can perform any task that is achievable through a sequence of primitive operations (e.g., typing or clicking) within an IDE; (ii) an agent has access to any tool implemented within the IDE since using a tool amounts to performing a sequence of primitive actions.
3.1 PwP environment
We represent the PwP environment as a partially observable Markov decision process (POMDP). We define the PwP POMDP $\langle S,A,O,T,R\rangle$ as:
  • $\mathbf{S}$ is the set of states describing the IDE and the operating system context, including open files, active editor panels, and cursor positions.
  • $\mathbf{A}$ is the action space, encompassing all possible keyboard and mouse events. The atomic actions in PwP are provided by the xdotool library (Sissel, ), which allows specifying all possible keyboard and mouse events in a simple syntax. The specific action space varies based on the agent setting, described in (§5).
  • $\mathbf{O}$ is the observation space. The observation space varies based on the agent setting, described in (§5).
  • $\mathbf{T}$ is the transition function. Actions like inserting a character typically lead to deterministic changes in the IDE state, whereas background processes can introduce stochasticity in timing and responses.
  • $\mathbf{R}$ is the reward function that measures performance on a given task. For instance, after the agent finishes editing code to fix a bug, the environment can run a test suite on the updated files to compute a reward.
Trajectories in PwP can thus resemble real-world development work: an agent can fix a bug in a repository, use a suggestion tool to help with writing code, or create documentation. The IDE and its operating system environment track changes, run tests and return reward signals.
Table 1: Comparison of PwP with existing environments across different dimensions. PwP uniquely combines comprehensive IDE tool support with full multimodal support, general action space, and execution-based evaluation, while maintaining software engineering specificity.
Multi- General Observation State Tools Execution-Based SWE
Environment modal Action Space Space Checkpointing Reward Specific
GAIA (Mialon et al., 2023) Text Limited
SWE-Bench (Jimenez et al., 2023) Text Limited
SWE-Bench-MM (Yang et al., 2024b) Text, Image Text Limited
OpenHands (Wang et al., 2024b) Text, Image Tool Output + Browser SWE
TheAgentCompany (Xu et al., 2024) Text, Image Tool Output + Browser SWE
WEBSHOP (Yao et al., 2023) Text Browser
WEBARENA (Zhou et al., 2024) Text Browser
VWEBARENA (Koh et al., 2024a) Text, Image Screen Browser
BrowserGym (Chezelles et al., 2024) Text, Image Tool Output + Browser Browser
OSWORLD (Xie et al., 2024) Text, Image Screen OS
WindowsAgentArena (Bonatti et al., 2024) Text, Image Screen OS
PwP (Ours) Text, Image, Video, Audio Screen All IDE Tools
3.2 Key Features of Programming with Pixels
Expressive observation and action space.
A typical approach to building agents is to engineer a set of high-level actions for operations like “open file” or “list symbols in file”, and then engineer an environment that supports each action (Xia et al., 2024; Yang et al., 2024b; Wang et al., 2024b). Furthermore, agents receive textual outputs that are manually reformatted for each action. The key difficulty is that such engineering does not scale to a large number of actions or to the full range of software engineering tasks, and the agent may be specialized to the observation and action space. In contrast, PwP preserves standard screen-based user interaction. The agent can navigate IDE menus visually, move the cursor, and press keys. This makes the environment expressive: an agent can achieve any task that can be achieved through primitive actions in an IDE.
Full Spectrum of Developer Tools.
A modern IDE offers debuggers, linters, version control, refactoring utilities, integrated terminals, and many extensions. In particular, PwP is developed on top of VSCode, which has a rich set of built-in functionalities and extensions. Implementing each of these into an agent or environment would require significant human effort. However, PwP provides these capabilities out-of-the-box within a single environment. Agents can set breakpoints, execute code, use language-specific extensions, review error messages, or run tests in a consistent interface.
Multimodality and Language Agnosticism.
Because the IDE supports numerous programming languages through extensions and built-in modules, PwP naturally covers tasks across Python, Java, JavaScript, Lean, and more without requiring separate integrations. For instance, agents can use the same debugger interface for all languages, or use pre-existing linters provided by IDE extensions. Beyond screenshots, the environment provides video streams and audio, though we leave exploring these for future work.
Rich Feedback and State Access.
PwP can evaluate performance immediately using testing frameworks or compilation checks. For instance, if the agent modifies a file, the IDE can automatically trigger a build, update diagnostics, or run tests, generating real-time feedback. In cases requiring deeper inspection—such as verifying that a bug is truly fixed—the environment can reveal file-system changes, process states, or test results.
1bench = PwPBench(dataset=’swebench’)
2# Replace with any dataset from PwP-Bench
3dataset = bench.get_dataset()
4
5# Set up environment and get initial observation
6env = bench.get_env(dataset[0])
7observation: PIL.Image = env.get_observation()[’screenshot’]
8
9# Generate and execute action
10action = agent.get_action(observation)
11print(action)
12# Output: xdotool mousemove 1000 1200
13# click 1 && xdotool type ’hello world’
14observation, info = env.step(action)
15
16env.render()
17
18# Environment control
19env.pause()
20env.resume()
21
22# Get reward and reset
23is_success = env.get_reward()
24env.reset()
Figure 2: Example demonstrating interaction with PwP environment, including keyboard/mouse actions, checkpointing, and state management. The code shows basic initialization, action execution, environment control, and reward handling.
Future Adaptability.
As agents continue to evolve, PwP provides a unified environment for incorporating new benchmarks and training. For example, PwP is amenable to reinforcement learning due to its use of the standard gymnasium (Towers et al., 2024) interface and its checkpointing functionality. We show an example of how to interact with PwP in Figure 2. Checkpointing also allows for backtracking in search-based methods (Koh et al., 2024b; Putta et al., 2024). As software engineering practices progress and new IDE tools emerge, PwP incorporates them without additional engineering overhead. Further, as agents become more ubiquitous, it is imperative to evaluate their capabilities in pair-programming scenarios with human developers. PwP supports concurrent user-agent interaction, potentially enabling new kinds of pair programming or real-time collaboration studies. Finally, adding new tasks requires only minimal modification to configuration and evaluation files.
3.3 Infrastructure and Implementation
PwP is deployed in a secure sandboxed environment. In particular, we run a modified version of Visual Studio Code (VSCode) and a minimal operating system inside a Docker container, ensuring a secure and isolated environment. We chose VSCode for its extensive language support, rich ecosystem of extensions, widespread adoption in the developer community, and open-source nature that enables customization and modification of its core functionality. Each container instance maintains its own file system and processes, preventing interference between experiments, facilitates reproducibility, and ensuring parallelization of evaluation. We further provide the ability to checkpoint the environment state, which is especially useful for backtracking in search algorithms or while training RL agents.
The environment interfaces with VSCode through multiple channels: 1.) A controller that manages Docker container lifecycle and configuration, 2.) A port-forwarding system for real-time screen and video capture, 3.) A modified VSCode codebase that exposes DOM state information, and 4) The VSCode Extension API for accessing fine-grained IDE state. This multi-channel approach enables both high-level environment control and detailed state observation.
Screen capture is handled via ImageMagick for static screenshots and ffmpeg for streaming video output. These tools were selected for their low latency and ability to handle various screen resolutions and color depths. For actions, a lightweight controller executes xdotool commands within the container, which in turn simulates keyboard and mouse events on the IDE. Agents can thus insert code, open new files, or navigate menus using the same actions that a human developer would.
As shown in 2, a Python API is provided for interaction, following a style similar to common reinforcement learning libraries such as gymnasium (Towers et al., 2024). The API abstracts away the complexity of container management, benchmark management, and handling observations and actions, allowing researchers to focus on agent development. Users can query the environment for the latest screenshot, issue an xdotool command, and receive updated states or rewards. The environment’s container configuration is flexible, allowing for software installations, customizable CPU/memory limits, and display settings (e.g., resolution). This versatility is crucial for large-scale evaluation, especially when tasks vary in complexity and resource needs.
4 PwP-Bench
We introduce PwP-Bench, a benchmark comprising 15 diverse software engineering tasks that span 8 programming languages and multiple modalities. Each task provides agents with access to the complete suite of tools available in the PwP environment. The purpose of PwP-Bench is to evaluate how well agents handle a broad range of software engineering (SWE) activities, thereby testing the generality of their code generation and SWE capabilities.
Tasks.
PwP-Bench contains 5400 instances covering 15 tasks, sourced from 13 existing code-generation datasets and 2 newly created by us. These tasks are designed to be representative of the breadth of software engineering—including tasks beyond conventional code generation to capture real-world complexities—and can be expanded as models excel in newer tasks. We followed three guiding principles: (1) tasks should require significant interaction with various SWE tools, (2) each task should necessitate multiple steps to complete, and (3) the overall benchmark should span multiple programming languages and modalities. Based on these principles, we collected diverse tasks and grouped them into four categories:
  • Code Generation and Editing: Evaluates the ability of agents to generate and edit code. This category includes datasets such as HumanEval for code completion, SWE-Bench (Jimenez et al., 2023) and SWE-Bench-Java (Zan et al., 2024) for resolving pull requests, DSBench (Jing et al., 2024) for data science tasks, and Res-Q (LaBash et al., 2024) or CanITEdit (Cassano et al., 2024) for code editing. Each dataset benefits from different tools. For example, SWE-Bench can take advantage of debuggers and linters, while DSBench may leverage an IPython kernel and tools for analyzing large data files. Code editing tasks can leverage refactoring utilities and repository searches, covering varied input-output formats and end goals.
  • Multimodal Code Synthesis: Involves creating code based on input images or other visual data. Examples include Design2Code (Si et al., 2024) for UI development, Chart2Mimic (Shi et al., 2024) for generating Python code from chart images, SWE-Bench-MM (Yang et al., 2024b) for multimodal code editing, and DSBench tasks that rely on images or PDF documents during data analysis.
  • Domain-Specific Programming: Focuses on specialized fields such as ethical hacking (CTF) (Yang et al., 2023a) and interactive theorem proving (miniCTX) (Hu et al., 2024). These tasks demand significant interactivity with IDE components. For example, theorem proving requires continuously inspecting states via the IDE, while CTF tasks often involve analyzing images, running executables, or installing VSCode extensions (e.g., hexcode readers).
  • IDE-Specific and General SWE Tasks: Recognizing that code generation is only one aspect of software engineering, we introduce two novel task sets to evaluate broader SWE skills. The first, IDE Configuration, evaluates an agent’s ability to modify IDE settings—such as themes, extension installations, and preferences—that are critical for effective tool use in a complex environment. The second, which we term General-SWE, targets non-code activities such as profiling, designing UI mockups, managing Kanban boards, and project refactoring. These tasks capture essential operational skills typically required by human developers but largely absent from conventional code generation benchmarks.
Figure 3 shows the distribution of tasks across all categories. An agent that succeeds across these tasks demonstrates strong potential for automating a wide range of software engineering activities. For instance, PwP-Bench covers Python, Java, JavaScript, HTML, CSS, Bash, SQL, and Lean, and requires agents to work with text, images, data files, and other data types. Furthermore, effective interaction with IDE tools is essential.
Figure 3: Distribution of tasks in PWP-Bench across four main categories: Code Generation and Editing, Multimodal Code Synthesis, Domain-Specific Programming, and General SWE Tasks. The inner ring shows the main categories while the outer ring shows specific datasets and tasks within each category.
Table 2: Comparison of existing software engineering benchmarks. PwP-Bench provides the largest dataset (5400 instances) and uniquely covers all aspects: multiple languages and modalities, real IDE interaction, interactive coding, and both code generation and general software engineering tasks.
#Instances Multiple Multiple Real IDE Interactive Non-Code Code-Generation
Benchmark Languages Modalities Env Coding SWE Tasks SWE Tasks
SWE-Bench (Jimenez et al., 2023) 2K
SWE-Bench-MM (Yang et al., 2024b) $\leq 1K$
LiveCodeBench (Jain et al., 2024) $\leq 1K$
Aider Polyglot (Aider, 2024) $\leq 1K$
TheAgentCompany (Xu et al., 2024) $\leq 1K$
VisualWebArena (Koh et al., 2024a) $\leq 1K$
OSWORLD (Xie et al., 2024) $\leq 1K$
WindowsAgentArena (Bonatti et al., 2024) $\leq 1K$
PwP-Bench (Ours) 5.4K
Benchmarking Design and Task Setup.
Every task is evaluated within the PwP environment. Unlike traditional benchmarks that provide structured, well-formatted context (for example, supplying all relevant schemas in text-to-SQL), PwP-Bench presents agents with an IDE containing a codebase rich in information. Specifically, an agent is provided with an initial environment state $S_{i}$ and an instruction $I$. The agent’s goal is to update the codebase to satisfy $I$ and transition to a final state $S_{f}$. Only $S_{f}$ is evaluated, using execution-based criteria (e.g., running unit tests). This setup requires agents to autonomously discover and extract relevant information from files, directories, and other resources—mirroring the challenges faced in real-world software development.
Many tasks in PwP-Bench require extensive multi-turn interactions and can be time-consuming. To support large-scale evaluations, PwP enables parallelized testing in a sandboxed environment, ensuring both security and reproducibility. Tasks can also be configured to restrict or partially allow internet access based on experimental needs.
Furthermore, as software engineering tasks evolve with advancements in model capabilities, our benchmark is designed to grow over time. New tasks can be incorporated by creating simple setup scripts that define the IDE’s initial state and evaluation logic, ensuring that PwP-Bench remains modular and adaptable.
PwP-Bench-Lite.
Because PwP-Bench contains more than 5400 instances in total, running a full evaluation can be computationally expensive. To address this, we also provide PwP-Bench-Lite—a smaller subset of 300 instances, made up of 20 random samples per task. This subset preserves the overall difficulty and distribution while ensuring equal representation for each task, thereby making rapid experimentation more accessible.
5 Agents in Programming with Pixels
The primary objective of Programming with Pixels is to enable general-purpose SWE agents. We evaluate state-of-the-art agents based on vision-language models in our environment. Each agent operates in a turn-based manner, receiving a screenshot each turn and returning an action (keyboard or mouse action) to progress toward the goal. This design is the same as used in previous works (Xie et al., 2024; Koh et al., 2024a) and we refer them to as computer-use agents. In practice, most vision-language models struggle with raw image inputs. To mitigate this, we incorporate Set-of-Marks (SoM) (Yang et al., 2023b), in which the agent receives both the raw image and a parse of available interface element (e.g., buttons, text fields). The agent then interacts with element IDs instead of raw pixel coordinates.
We evaluate two categories of computer-use agents. The first category only outputs keyboard and mouse clicks. That is, the agent uses the following observation and action space:
  • $\mathbf{O}$: a screenshot and set-of-marks annotations.
  • $\mathbf{A}$: keyboard and mouse clicks with set-of-marks.
The second category of computer-use agents has access to file and bash commands supplied by the environment through an API. Furthermore, a screenshot is received only when the screenshot action is called, instead of receiving it every turn. These actions are provided in PwP in a similar design principle as Anthropic computer-use (Anthropic, 2024) which consists of file operations such as ‘read file’, ‘create file’, and ‘string replace’. In summary, these agents use the following observation and action space:
  • O: a screenshot and set-of-marks annotations and text output from tools if used.
  • A: keyboard, mouse, and file and bash operation actions.
Finally, in Analysis 6.2, we create a tool-based agent design compatible with PwP. Specifically, we provide agents with domain-specific API calls that represent high-level actions (such as getting file structure) instead of UI interactions. Each high-level action is implemented as a sequence of low-level actions executed in PwP. This setting lets us test current state-of-the-art SWE agent patterns (e.g., Wang et al. (2024b)) within the PwP environment.
6 Experiments
Table 3: Performance Evaluation of Different Agents on PwP-Bench by Task Categories
Inputs Outputs Model Code Generation Multimodal Domain-Specific General Overall
& Editing Code Generation Code Generation SWE Tasks Avg
Screenshot + SoM Keyboard + Mouse GPT-4o 0.8% 12.4% 1.7% 10.0% 5.3%
GPT-4o-mini 0.8% 3.7% 0.0% 2.5% 1.7%
Gemini-Flash 0.0% 4.3% 0.0% 0.0% 1.2%
Gemini-Pro 2.5% 5.7% 0.0% 7.5% 3.%
Claude-Sonnet 10.0% 8.2% 5% 20.0% 10.2%
Screenshot + SoM + Tool Output Keyboard + Mouse + Tool Call (File, Bash) GPT-4o 37.0% 41.9% 28.3% 5.0% 32.3%
GPT-4o-mini 23.6% 17.5% 15.0% 5.0% 17.8%
Gemini-Flash 9.5% 11.7% 8.3% 2.5% 8.9%
Gemini-Pro 30.9% 16.7% 3.3% 5.0% 18.1%
Claude-Sonnet 52.1% 55.1% 43.3% 20.0% 46.8%
Figure 4: Example of Successful Use of Live Preview Tool in the UI Replication Task The agent successfully uses the live preview tool in the VSCode browser to compare the UI design it made versus the reference design.
Figure 5: Example of Successful Use of Tool in the Chart Generation Task The agent can compare the generated chart with the reference chart side by side and refine its code accordingly.
Experimental setup.
We evaluate two categories of baseline agents as described in Section 5. For each configuration, we test five state-of-the-art vision-language models (VLMs): Gemini-Flash-1.5, Gemini-Pro-1.5, GPT-4o, GPT-4o-mini, and Claude-3.5 Sonnet. With the exception of Claude—which is natively trained for UI interaction—the remaining models are provided with SoM.
At each timestep, an agent receives an observation (with the observation space determined by its category) and returns an action. The complete history of observations and actions is incorporated into the model’s context. For each task instance, the maximum number of iterations is capped at 20 steps; if the agent either exhausts these steps or issues a stop command, the environment’s final state is evaluated using task-specific metrics (see Appendix B for full details). Notably, the agent design remains the same throughout all tasks. Due to computational and budget constraints, we evaluate on PwP-Bench-Lite, which has 300 task instances.
6.1 Results
Figure 6: Example of the Claude Computer-Use agent successfully using multiple IDE tools to complete a repository level code-editing task.
Figure 7: Example of Agent Hallucinating Screen Contents While the agent correctly mentions, search results are displayed (green text), it hallucinates an input field containing “disable import error” (red)
Figure 8: Example of wrong mouse click by Claude-Computer Use Agent The agent attempted to click Settings icon but clicked at the wrong location.
Table 3 summarizes performance across different agent architectures and base models over the four categories of PwP-Bench. As seen in the top half of the table, when using only primitive keyboard and mouse actions most agents have poor performance, with a maximum overall average of 10.2%. We attribute this poor performance primarily to limited visual grounding and an inability to interact effectively with the IDE—particularly for file editing and tool usage (see Section 6.2 for further analysis).
In contrast, when agents are granted access to file editing and bash operations through API calls rather than relying solely on UI interactions, we observe consistent improvements across all categories, with maximum average accuracy reaching 46.8%. Among the evaluated models, the Claude computer-use agent performs best, likely because it is specifically trained for UI interactions. We found that this agent is able to leverage basic IDE tools—such as HTML live preview, chart visualization, and file navigation—to boost performance on tasks requiring visual understanding and IDE navigation. Importantly, we find for the first time that a single computer-use agent can achieve performance comparable to and often surpassing (See Appendix C) state-of-the-art methods across a wide variety of software engineering (SWE) tasks—encompassing multiple languages, modalities, and domains—while operating within a single unified environment and interface, and no specific hand-crafted tools, unlike existing SoTA methods.
Nonetheless, as detailed in Section 6.2, the models struggle to fully take advantage of the IDE tooling. This is evidenced by the poor performance on the ‘General SWE’ dataset, where tasks are often as simple as editing IDE settings and often require fewer than four clicks to complete. As we show in Section 6.2, these tasks become simpler if the models could use the IDE tooling more effectively.
Overall, while the results point toward a promising direction for developing general computer-use SWE agents, significant improvements are still needed in visual grounding, tool usage, and planning. We analyze these next.
6.2 Analysis
Agents Demonstrate Poor Visual Grounding Capabilities.
Our qualitative analysis across multiple VLMs on PwP-Bench reveals significant limitations in visual understanding—even for basic IDE interactions. We identify two primary failure modes. First, models frequently fail to correctly identify the UI elements intended for interaction, as demonstrated in Figure 78. In agents using set-of-marks (SoM), this issue manifests as incorrect element selection, while without SoM it leads to inaccurate mouse positioning. Second, models struggle to comprehend the current UI state. As shown in Figure 910, they consistently fail to recognize highlighted elements, cannot detect linter errors indicated by wavy underlines (Figure 11), and often confuse active panels—resulting, for example, in typing into search bars rather than file editors (Figure 9). While similar issues have been documented in web and OS domains (Koh et al., 2024a; Xie et al., 2024), these limitations were primarily observed in models without UI-specific training. However, our work, shows even models explicitly trained for UI interaction (Anthropic, 2024), including Claude-Computer Use, exhibit these issues in PwP—likely due to the increased complexity of the IDE interface.
Agents Fail to Edit Files. File editing is a basic capability required in most SWE tasks. However, we find that the deficiencies in visual grounding significantly impact the file editing capabilities of current agents that use basic actions (clicking and typing). For example, even when provided with cursor location information in textual form, these models struggle to interpret such data amid complex UI elements. Models fine-tuned for UI interactions still commit basic editing errors—such as incorrect indentation and text misplacement—and are unable to recover from these errors (see Appendix for examples). We speculate these limitations could stem from two factors: (i) model overfitting to user interfaces in their training domains, or (ii) the increased complexity of the PwP IDE interface, which contains substantially more interactable elements than typical web or OS environments. Addressing these limitations represents an important direction for future work. Although direct file access via tool operations is available, UI-based editing confers unique advantages for tasks such as editing Jupyter notebooks, comparing changes, or modifying specific sections of large files. These results underscore two limitations: (i) current VLMs are challenged by complex UI interactions beyond simple web/OS interfaces (Xie et al., 2024; Koh et al., 2024a), and (ii) the inability to effectively perform UI-based editing prevents agents from leveraging valuable IDE features that could have improved their performance.
Figure 9: Example of Agent Misidentifying UI Elements The agent fails to identify the correct input field, typing ‘50’ into the settings search bar instead of the word wrap column setting field (red arrow).
Figure 10: Example of Agent Misidentifying Active Panel The agent fails to recognize the active editor panel, incorrectly typing into the search bar (red arrow) instead of the file editor.
Claude Computer-Use Agent Demonstrates Basic IDE Tool Proficiency.
Our analysis shows that the Claude Computer-Use agent successfully uses fundamental IDE functionalities, including file explorer navigation, file editing, search, browser-based live preview, and image generation and visualization capabilities. Figure 4 demonstrates the agent’s effective use of browser tools in UI replication tasks. Similarly, Figure 6 illustrates the agent’s ability to coordinate multiple tools while editing specific lines in a repository, relying solely on screenshot observations and primitive keyboard/mouse actions.
Second, we hypothesize that the agent has additional abilities to use tools that can be uncovered through prompting or fine-tuning. To investigate this, we considered the symbol renaming task in our ‘General-SWE’ benchmark, in which Claude initially achieves 0% accuracy when attempting the task. However, when explicitly instructed to use the renaming tool, its accuracy improves to 50% (see Appendix C).
Agents Struggle to Use IDE Functionality.
Despite these successes with basic IDE functionality, computer-use agents demonstrate significant limitations when interfacing with more complicated IDE tools. We observe no successful instances of debugger usage or symbol listing operations. Models without specific UI interaction training struggle even with fundamental tools such as HTML live preview, image visualization, and graph generation capabilities. While Claude demonstrates competency with basic tools, it fails to effectively use advanced tools like profilers or debuggers.
To further evaluate these capabilities, we developed the ‘General-SWE’ dataset, which focuses on software engineering activities (e.g., profiling, refactoring, debugging) that can be completed without direct code modification. Although these tasks sometimes require only 4-5 steps when using appropriate IDE tools, agents achieve minimal performance, highlighting substantial room for improvement.
Training Models to Use IDE Tools Better Would Improve Performance.
While a single computer-use agent design can perform well across a wide variety of tasks, our results indicate that these models do not fully exploit domain-specific tools. As an indication of the potential for performance gains if the agent was able to effectively use the IDE, we perform an “assisted” experiment.
For the assisted experiment, we manually engineer a set of API calls that are useful for the tasks. For example, in Design2Code, the assisted agents have an API call for a live HTML preview, while for SweBench it has API calls for retrieving repository structure and symbol outlines. Importantly, each API call is achievable using basic operations in the IDE, meaning that in principle an agent could learn to perform it. To ensure that each API call is achievable in the IDE, we implement each API call by executing a fixed sequence of low-level IDE actions, with the details abstracted away from the agent. A complete list of tools available in different environments is in Table 7 (see Appendix C).
Table 4 compares the performance of these assisted agents with that of standard computer-use agents across four datasets for which we manually created tools. The assisted agents achieved up to a 13.3% improvement in average scores relative to the non-assisted agents. This suggests that training agents to explore and use the built-in IDE functionality would yield performance gains. It also suggests that in the near term, we can get performance gains by introducing hand-engineered tools into the computer-use agent and incorporating existing agent designs in our PwP environment.
Table 4: Comparison of Different Agent Types Across Selected Tasks
SWE-Bench Design2Code Chartmimic BIRD (T2 SQL)
Primitive Agents 0% 23.5% 2.7% 0%
Computer Use 15% 48.1% 25.3% 7%
Assisted 19% 79.5% 61.6% 17%
Figure 11: Example of Agent Missing Visual Error Indicators The agent fails to recognize linter error indicators (wavy underlines).
Figure 12: Example of Agent’s Inability to Perform File Editing The agent incorrectly positions new content in the file editor.
Agents Are Incapable of Recovering from Errors. Next, we find that current agents show limited error recovery capabilities. When an action fails to execute correctly, models tend to persistently repeat the same failed action without exploring alternatives. Similarly, if an agent selects an incorrect action, it continues along an erroneous solution path without recognizing or correcting the mistake. In an experiment designed to probe this behavior, we deliberately suppressed one of the model’s actions. Despite the environment’s screenshot clearly showing an unchanged state, the models proceeded with their planned action sequence as though the suppressed action had succeeded. This behavior suggests a heavy reliance on memorized action sequences rather than dynamic responses to visual feedback, resulting in exponentially increasing errors and poor performance.
7 Conclusion
In this work, we introduce PwP, a unified environment that challenges the prevailing tool-based paradigm by enabling direct interaction with IDEs through basic computer-use actions like typing and clicking. This approach allows for a wide range of tasks to be modeled without language- or domain-specific modifications. Second, we introduce PwP-Bench, a unification of existing SWE datasets evaluated in PwP. We find that general-purpose computer-use agents can approach or outperform previous state-of-the-art results, without any task-specific modifications. This suggests that the dominant paradigm of building specialized text-based tools for SWE agents may be superseded by end-to-end computer-use agents. However, our analysis reveals that state-of-the-art agents are still incapable of using the extensive set of tools available in PwP, and could perform better if they could use them. Our work opens up an exciting new direction of development of computer-use agents for SWE tasks, an important step towards reaching truly general-purpose SWE agents. Finally, PwP provides a common platform that supports existing agents and benchmarks, making it a foundation for future research on SWE agents.
Acknowledgments.
We thank Daniel Fried, Graham Neubig, and Zora Wang, Saujas Vaduguru, Atharva Naik, Riyaz Ahuja, Weihua Du for helpful feedback, Google and Open AI credit programs, and Convergent Research.
Appendix A Programming with Pixels (PwP) Environment
A.1 Tools
Previous methods have proposed use of various hand-engineered tools. However, as shown in Table 5, PwP natively supports all these tools.
Table 5: Comparison of Hand-engineered Tools across Methods versus PwP. PwP natively supports all tools.
Method Hand-engineered Tools Supported in PwP
Agentless (Xia et al., 2024) File Edit, Repository Structure, File Structure
CodeAct (Wang et al., 2024a) File Edit, IPython, Bash
SWE-agent (Yang et al., 2024a) Search File, Search Text, File Edit
EnIGMA (Abramovich et al., 2024) SWE-agent Tools + Debugger, Terminal, Connection Tool
swebench-mm (Yang et al., 2024b) SWE-agent Tools + View Webpage, Screenshot, Open Image
A.2 Comparison with Other Environments
In Table 1, we compare PwP with existing environments across multiple dimensions. We focus on environments designed for software engineering agents as well as those developed for web agents and operating system interaction. Since many prior code agent efforts do not explicitly separate their agent and environment architectures, we evaluate their environments based on their published interfaces (e.g., SWE-Bench, SWE-Bench-MM). We evaluate environments along the following dimensions:
  • Multi-modal support: The ability to process diverse input modalities including text, images, video, and audio
  • General action space: Whether the environment provides a general-purpose interface for tool interaction. For instance, OSWorld, WindowsAgentArena, and BrowserGym enable keyboard/mouse interactions through visual inputs, creating a general action and observation space. In contrast, environments like SWE-Bench and OpenHands restrict interactions to predefined tool-specific APIs. While OpenHands does allow BrowserGym interface, which in turns allows for general action and observation space, the generality is limited to browser interactions and does not apply to other SWE-related tools.
  • Observation space: The range of environmental states and feedback accessible to an operating agent
  • State Checkpointing: Native support for preserving file system and process states, a capability unique to PwP
  • Tools: The set of available tools and capabilities accessible to agents
  • Execution-based evaluation: Use of runtime execution to verify the correctness of agent actions
  • SWE-specific: Whether the environment is purposefully designed for software engineering tasks
While some environments offer general action and observation spaces, they are not specifically designed for software engineering tasks. Programming with Pixels uniquely combines SWE-specific design with general-purpose interaction capabilities, enabling the development of general-purpose computer-use agents for software engineering.
Appendix B PwP-Bench
Metrics
We use individual metrics mentioned in the original datasets. When reporting results on PwP-Bench, we report marco average of all these metrics. In particular, 11/15 used Accuracy as their metric. However, due to complexity of dataset, these often goes beyond simple accuracy metrics and in some cases, the dataset is evaluated on multiple orthogonal metrics, instead of one. We detail, these metrics for each of the datasets.
  • SWT-Bench evaluates generated tests by the agent, and reports 6 different metrics: Applicability, Success Rate, F- X, F- P, P- P, and Coverage. We report the average of all 6 metrics.
  • ChartMimic evaluates generated code on various metrics such as accuracy of text, colors used, legend etc. We average all metrics similar to the original dataset.
  • Design2Code evaluates generated code on various metrics such as accuracy of text, position, clip score, etc. We average all metrics similar to the original dataset.
  • DSBench has two categories, one containing MCQ questions, while the other containing generating code for Kaggle Competitions. We use 10/10 instances from each category in PwP-Bench-Lite. While MCQ questions are evaluated using Accuracy, the code generation part is evaluated using linear normalization between the baseline score (of the competition) and the score of the winner of competition.
Appendix C Results
Table 6 presents comprehensive results for all agent designs across 15 datasets in PwP-Bench. For comparison, we include previously reported state-of-the-art results, which represent the best performance achieved on individual datasets through various specialized approaches including finetuned models, custom tool interfaces, specific pipelines, prompts, inference strategies, and verifiers.
It is important to note that direct comparisons on individual datasets may not provide a complete picture, as state-of-the-art results often utilize different model versions, computational budgets, and domain-specific tools. For example, the 55% performance on SWE-Bench was achieved using significantly longer agent trajectories compared to PwP’s 20-iteration limit. Given statistical variations and differing evaluation conditions, we emphasize evaluating agents across the full range of tasks rather than focusing on individual dataset performance, as this better aligns with our goal of developing general-purpose SWE agents. We make our best effort to include the latest publicly available results, however, there may be minor discrepancies.
Notably, on the Intercode dataset (Capture the Flag), the computer-use agent significantly outperforms the state-of-the-art method despite the latter (SWE-Agent-Enigma) employing numerous specialized tools for static analysis, dynamic analysis, and networking (See Table 5). This improvement is statistically significant (p-value = 0.014, McNemar’s test). While the agent did not rely on specific IDE tooling, this result highlights the advantages of a simple, general agent design and suggests the potential benefits of allowing agents to automatically identify optimal tools rather than imposing human priors on the system.
Table 6: Performance Evaluation of Different Models Across Task Categories. Leged: HE: HumanEval, SB: SWEBench, SJ: Swebench-Java, RQ: ResQ, CI: CaniteEdit, ST: SWTBench, DC: Design2Code, CM: ChartMimic, DS: DSBench, SM: Swebench-MM, IC: Intercode-CTF, BD: Bird SQL, MC: Minictx, VS: VSCode, GS: No-Code SWE Tasks. *Much more costly method, ** Estimated Value since only Pass@8 is reported.
Code Generation & Editing Multimodal Domain-Specific No-Code
Code Generation Code Generation SWE Tasks Overall
Model HE SB SJ RQ CI ST DC CM DS SM IC BD MC VS GS Avg
Computer-Use Agents   (Screenshot + SoM)
GPT-4o 5% 0.0% 0.0% 0.0% 0.0% 0.0% 48.7% 0.7% 0.0% 0.0% 5.0% 0.0% 0.0% 20.0% 0.0% 5.3%
GPT-4o-mini 0.0% 0.0% 0.0% 0.0% 5.0% 0.0% 14.8% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 5.0% 0.0% 1.7%
Gemini-Flash 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 15.2% 2.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 1.2%
Gemini-Pro 10.0% 0.0% 0.0% 0.0% 5.0% 0.0% 14.5% 8.1% 0.0% 0.0% 0.0% 0.0% 0.0% 15.0% 0.0% 3.5%
Claude-Sonnet 20.0% 0.0% 0.0% 15.0% 25.0% 4.2% 18.1% 0.0% 5.0% 10.0% 15.0% 0.0% 0.0% 35.0% 5.0% 10.2%
Computer-Use Agents   (Screenshot + SoM + File/Bash Operations)
GPT-4o 85% 25% 15% 30% 50% 17.0% 70.2% 65.5% 11.9% 20% 70% 10% 5% 10% 0.0% 32.3%
GPT-4o-mini 60% 10% 5% 20% 30% 16.7% 41.3% 5.5% 8.4% 15% 40% 5% 0% 10.0% 0.0% 17.8%
Gemini-Flash 0 5% 5% 15% 15% 17.1% 19.9% 13.5% 3.2% 10% 25% 0% 0% 5% 0.0% 8.9%
Gemini-Pro 85% 10% 15% 15% 40.0% 20.2% 25.6% 24.7% 1.6% 15% 5% 5% 0% 10% 0.0% 18.1%
Claude-Sonnet 95% 25% 35% 55% 65% 37.4% 83.4% 71.2% 55.7% 10% 100% 15% 15% 35% 5.0% 46.8%
Previous State of the Art Reported*
98.8% 55%* 9.9% 58% 63.3% $37.1\%$ 90.2% 71.4% 39.4% 12.2% 72% 30.2% $\approx 25\%$** 35.0% 5.0% 46.8%*
Appendix D Additional Results
Table 7: Tools available in different environments. The table shows the various tools provided by different environments for assisted analysis. Common tools like file manipulation and bash operations are shared across environments, while specialized tools cater to specific tasks like web design and chart replication.
Category Tool Description
Common Tools bash Perform bash operations
file_edit Perform file manipulation operations
SWEBench search_repository Search the repository for a string in the entire repository
file_name_search Search for a file by its name
view_structure View the structure of the current directory
Design2Code view_html_preview Get a preview of the index.html page as rendered in the browser
view_original_image Get a screenshot of the html image for replication
zoom_in Zoom in on the current rendered html page
zoom_out Zoom out on the current rendered html page
ChartMimic view_python_preview Get a preview of the graph generated by python file
view_original_image Get a screenshot of the graph for replication
BIRD test_sql Test a SQL query against the database
get_relevant_schemas Get relevant descriptions of the relevant database tables
Training models to use IDE tools better would improve performance.
In Section 6.2, we demonstrate that models can achieve superior performance when effectively utilizing IDE tools. However, our analysis reveals two primary limitations in current models’ tool usage: (1) poor visual grounding and inability to handle complex tool interfaces, and (2) failure to prioritize IDE tool-based solutions over manual approaches.
To evaluate the second limitation specifically, we developed refactoring tasks within our ‘General-SWE’ dataset. These tasks require agents to rename symbols across a project repository—an operation that cannot be reliably accomplished through simple search-and-replace due to potential naming conflicts and contextual variations. The IDE provides a robust solution through its rename feature, which leverages the complete AST to ensure accurate symbol renaming across the codebase. This operation requires only pressing F2 on a symbol and entering the new name. In our evaluation, the Claude agent initially achieved 0% accuracy across four tasks when given no tool guidance. However, when explicitly prompted with ”You can utilize the rename feature in VSCode to perform this task,” its accuracy improved to 50%.
We observed similar patterns across other tasks designed to evaluate tool usage. For instance, tasks that could be efficiently solved using the debugger showed limited success. While agents could sometimes set breakpoints, their poor visual grounding prevented them from effectively interpreting the debugging interface—particularly in understanding the current execution state and paused line location. These findings suggest significant potential for improving agent performance through better training on IDE tool utilization.
Successful Use of Tools
We further show a couple of examples of successful tool use in Figure 4 5. However, we do note that while the agent is able to use the IDE tool through UI interaction, it still may not be able to make optimal use of it as shown in Figure 13.
Figure 13: Performance comparison of GPT-4 agent in Computer-Use and Assisted settings on the ChartMimic dataset. a) Image as seen by the Computer-Use agent. b) Replication in Computer-Use setting. c) Replication in Assisted setting. The Assisted agent demonstrates superior performance despite seeing the same image but in different context and state.

### v2

@@
-Programming with Pixels: Computer-Use Meets Software Engineering
+Programming with Pixels: Can Computer-Use Agents do Software Engineering?
 Abstract
-Recent advancements in software engineering (SWE) agents have largely followed a tool-based paradigm, where agents interact with hand-engineered tool APIs to perform specific tasks.
-While effective for specialized tasks, these methods fundamentally lack generalization, as they require predefined tools for each task and do not scale across programming languages and domains.
-We introduce Programming with Pixels (PwP), an agent environment that unifies software development tasks by enabling computer-use agents—agents that operate directly within an IDE through visual perception, typing, and clicking, rather than relying on predefined tool APIs.
-To systematically evaluate these agents, we propose PwP-Bench, a benchmark that unifies existing SWE benchmarks spanning tasks across multiple programming languages, modalities, and domains under a task-agnostic state and action space.
-Our experiments demonstrate that general-purpose computer-use agents can approach or even surpass specialized tool-based agents on a variety of SWE tasks without the need for hand-engineered tools.
-However, our analysis shows that current models suffer from limited visual grounding and fail to exploit many IDE tools that could simplify their tasks.
-When agents can directly access IDE tools, without visual interaction, they show significant performance improvements, highlighting the untapped potential of leveraging built-in IDE capabilities.
-Our results establish PwP as a scalable testbed for building and evaluating the next wave of software engineering agents.
-[^1] Carnegie Mellon University.
-We release code and data at programmingwithpixels.com.
-Machine Learning, ICML
+Computer-use agents (CUAs) hold the promise of performing a wide variety of general tasks, but current evaluations have primarily focused on simple scenarios.
+It therefore remains unclear whether such generalist agents can automate more sophisticated and specialized work such as software engineering (SWE).
+To investigate this, we introduce Programming with Pixels (PwP), the first comprehensive computer-use environment for software engineering, where agents visually control an IDE to perform diverse software engineering tasks.
+To enable holistic evaluation, we also introduce PwP-Bench, a benchmark of 15 existing and new software-engineering tasks spanning multiple modalities, programming languages, and skillsets.
+We perform an extensive evaluation of state-of-the-art open-weight and closed-weight CUAs and find that when interacting purely visually, they perform significantly worse than specialized coding agents.
+However, when the same CUAs are given direct access to just two APIs—file editing and bash operations—performance jumps, often reaching the levels of specialized agents despite having a task-agnostic design.
+Furthermore, when given access to additional IDE tools via text APIs, all models show further gains.
+Our analysis shows that current CUAs fall short mainly due to limited visual grounding and the inability to take full advantage of the rich environment, leaving clear room for future improvements.
+PwP establishes software engineering as a natural domain for benchmarking whether generalist computer-use agents can reach specialist-level performance on sophisticated tasks.
+[^1] Code and data released at https://programmingwithpixels.com
 1 Introduction
-Figure 1: Comparison between the traditional tool-based paradigm (left) and our proposed Programming with Pixels (PwP) framework (right) for software engineering (SWE) agents.
-Instead of relying on specialized hand-engineered tools, PwP enables agents to interact directly with an IDE through basic computer interactions and screen observations.
-The framework naturally allows agents to take advantage of existing IDE capabilities across multiple languages and tools.
-Furthermore, PwP-Bench is a comprehensive benchmark to evaluate agent performance across different SWE domains.
-Human software developers possess a remarkable ability to work across a wide range of programming tasks, seamlessly adapting to new languages, tools, and problem domains.
-Realizing a single, general-purpose agent with similar versatility is the overarching goal of many recent efforts in code generation and software engineering automation (Jiang et al., 2024; Jin et al., 2024; Wang et al., 2024b).
-However, most software engineering agents (SWE agents) still rely on a tool-based paradigm, where an agent takes actions using hand-engineered functions (e.g., search repository, run Python code) exposed through a text API (Yang et al., 2024a, b; Wang et al., 2024a, b).
-This fundamentally limits generalization, since tool-based agents can only perform tasks using the predefined actions.
-For example, an agent designed to manage GitHub pull requests lacks debugging abilities unless it is programmed into the agent’s API.
-Furthermore, the tool-based paradigm lacks scalability, as hand-engineering complex tools requires significant human effort and may not be bug-free.
-As a result, it remains unclear whether the tool-based paradigm scales to the diversity of software engineering tasks, which spans multiple languages, modalities, and task types.
-Our motivating hypothesis is that achieving general-purpose SWE agents requires a shift to computer-use agents (Anthropic, 2024) that interact with computers as humans do: by observing the screen, typing, and clicking.
-To this end, we recast agentic software engineering as interacting directly with an integrated development environment (IDE) by observing its visual state and using basic actions such as clicking and typing.
-This allows the agent to perform any task possible in an IDE and leverage all of the IDE’s tools—from debuggers to web browsers—without requiring specialized APIs.
-However, despite promising results in web navigation (Anthropic, 2024) and open-ended computer tasks (Xie et al., 2024), the ability of computer-use agents to perform software engineering remains underexplored and we lack a dedicated environment for software engineering.
-To close this gap, we introduce Programming with Pixels (PwP), the first software engineering agent environment aimed at general-purpose computer-use agents.
-The PwP environment is a VSCode-based IDE where agents perceive the screen and use primitive actions such as typing, pointing, and clicking.
-PwP fulfills two key properties.
-First, the environment is expressive, allowing agents to complete any software engineering task achievable in an IDE, without language- or domain-specific modifications.
-Second, agents naturally interact with any tools available in the IDE–including debuggers, linters, and code suggestions while handling diverse data types such as images, videos, and PDFs–through basic actions such as clicking and typing.
-This notion of tool use fundamentally differs from hand-engineered tool APIs, offering scalability and reducing the effort needed to hand-engineer tools for AI agents.
-Namely, PwP lets agents take advantage of the rich tools already accessible to humans in an IDE, rather than reinventing the wheel.
-Finally, computer-use agents reduce the need for complex tool pipelines (e.g., tool-specific prompts), opening up a simplified approach to general-purpose SWE agents.
-To further evaluate agents developed for computer-use we construct PwP-Bench, a unified benchmark of 15 tasks spanning a variety of software engineering activities, including code generation, pull request resolution, UI development, and DevOps workflows.
-We show for the first time that general-purpose computer-use agents achieve non-trivial performance on a wide variety of SWE tasks, often approaching or exceeding state-of-the-art tool-based agents.
-However, our analysis reveals substantial opportunities for future work.
+Computer-use agents (CUAs) hold the promise of automating a wide range of economically valuable tasks by acting through primitive actions such as clicking, typing, and observing digital screens, potentially obviating the need for specialized AI agent action interfaces (Anthropic, 2024; OpenAI, 2025; Yang et al., 2024a).
+However, current evaluations have primarily focused on simple tasks such as web navigation (Koh et al., 2024), basic document editing, or tweaking settings in operating systems (Xie et al., 2024; Bonatti et al., 2024).
+Therefore, it remains unclear whether current generalist computer-use agents can automate more sophisticated and specialized tasks such as software engineering.
+In this work, we specifically study how well the current generation of computer-use agents can do software engineering and identify their key limitations.
+The choice of using software engineering as the test domain is motivated by two primary reasons.
+First, software engineering represents an economically important and practically challenging task.
+Second, the field of AI software-engineering agents (SWE agents) has produced numerous specialized agents that use hand-engineered APIs for specific operations (Yang et al., 2024a; Wang et al., 2024b; Xia et al., 2024), providing strong baselines for comparison.
+These agents use custom functions such as file editing, code search, and repository management, with each tool requiring significant engineering effort and domain expertise.
+For instance, SWE-agent (Yang et al., 2024a) uses language-specific parsers and editing commands, while Agentless (Xia et al., 2024) relies on Python-specific abstract syntax trees.
+This specialization has yielded strong performance, but it raises a fundamental question: can general-purpose computer-use agents match specialized agents in complex domains like software engineering?
+Figure 1: Programming with Pixels is an environment for computer-use agents, where they interact with a VSCode IDE through keyboard and mouse actions while observing the screen.
+The framework supports multiple programming languages, tests interactions with multiple IDE features, modalities (eg: text, images, data files).
+PwP-Bench evaluates agents across 15 diverse software engineering tasks such as code generation, UI generation, Data Science.
+To investigate this question, we introduce Programming with Pixels (PwP), the first environment for systematically evaluating computer-use agents on software engineering tasks.
+The PwP environment provides a VSCode-based IDE where agents perceive the screen and use primitive actions such as typing and clicking to perform a variety of SWE tasks.
+This design enables two critical properties for fair evaluation.
+First, the environment is expressive, allowing agents to complete any software engineering task achievable in an IDE without language- or domain-specific modifications.
+Second, agents can access all IDE tools—debuggers, linters, code suggestions—through the same visual interface available to human developers or specialized SWE agents.
+Hence, PwP provides a general-purpose, realistic software engineering environment for testing computer-use agents.
+To evaluate computer-use agents, we construct PwP-Bench, a benchmark of 15 tasks spanning different tasks such as code generation, pull request resolution, UI development, and data science across multiple programming languages and modalities.
+The benchmark represents a unification of 13 existing SWE tasks ported for evaluating computer-use agents, and 2 additional tasks developed by us.
+Our evaluation of state-of-the-art computer-use agents reveals that when restricted to pure visual interaction, these agents achieve only 22.9% average accuracy, significantly underperforming specialized coding agents.
+However, when augmented with just two basic text APIs—file editing and bash operations—the same agents achieve 50.7% accuracy, often approaching specialized agent performance despite their task-agnostic design.
+Furthermore, our analysis reveals substantial opportunities for future work.
 First, even state-of-the-art computer-use agents suffer from visual grounding issues.
-Second, we show that current agents lack the ability to use many of the tools available in the IDE, including ones that could make their tasks trivial.
-This suggests that training agents to explore and use the tools available in the IDE is a fruitful future direction.
-Finally, we find that only one model, Claude (Anthropic, 2024), performs well, highlighting the need for further research into training and improving computer-use agents.
+Second, we show that current computer-use agents lack the ability to use many of the tools available in the IDE, including ones that could make their tasks trivial.
+This suggests that training computer-use agents to explore and leverage the functionality present in their computer environment is a fruitful future direction.
+Overall, our results highlight software engineering as a realistic and challenging benchmark for evaluating and improving computer-use agents.
 In summary, our contributions are as follows.
 First, we introduce Programming with Pixels (PwP), the first software engineering-focused environment for evaluating computer-use agents.
-Second, we introduce PwP-Bench, a benchmark spanning 15 diverse SWE domains, allowing for systematic comparison of software engineering agents.
-Third, we demonstrate, for the first time, that computer-use agents can perform a wide variety of software engineering tasks without additional hand-engineering of their action or observation spaces.
-Fourth, we analyze the limitations of current computer-use agents, identifying the need for models that better leverage IDE tooling and highlighting agent training as a key future direction.
-Finally, both existing agents and benchmarks can be easily incorporated into our unified environment, positioning PwP to serve as a common platform for developing future SWE agents.
-Overall, PwP challenges the prevailing tool-based paradigm for SWE agents and provides a platform for developing more general agents that interact directly with IDEs.
+Second, we introduce PwP-Bench, a benchmark spanning 15 diverse SWE domains, allowing for systematic comparison of computer-use agents.
+Third, through extensive evaluation, we highlight the limitations of current computer-use agents, identifying the need for models that have better visual grounding and that better take advantage of their environment as key future directions.
+Finally, we open-source our environment and benchmark, allowing it to serve as an open platform for evaluating and improving agents on software engineering tasks.
 2 Related Work
-2.1 Task-specific SWE benchmarks
-Early neural code generation approaches were typically evaluated on fixed input-output pairs—for example, generating code from docstrings (Chen et al., 2021) or from general textual descriptions (Austin et al., 2021).
-Subsequent benchmarks extended these evaluations to interactive settings, such as resolving GitHub pull requests or writing unit tests for real-world code repositories (Jimenez et al., 2023; Zan et al., 2024; Mündler et al., 2025).
-More recently, efforts have broadened the scope of code generation to include multimodal tasks, where vision models must interpret images to generate correct code or edits (Si et al., 2024; Shi et al., 2024; Jing et al., 2024; Yang et al., 2024b).
-However, each of these benchmarks is confined to specific languages, modalities, or task types.
-In contrast, our proposed PwP-Bench unifies these diverse evaluations into a single framework, encompassing multimodal and multilingual challenges that require interaction with a broad suite of IDE tools.
-Using this unified approach we reproduce the performance of established benchmarks and encourage the development of general-purpose agents capable of handling a variety of new software engineering tasks.
-We further compare our work with previous efforts in Tables 1 and 2.
-2.2 Software Engineering (SWE) Agents
-Recent work has explored “code agents” that move beyond single-step neural code generation toward interactive methods, where intermediate feedback from tools informs subsequent actions.
-However, many of these approaches specialize in particular tools or programming languages (Jin et al., 2024; Yang et al., 2024b), limiting their broader applicability.
-For example, Agentless (Xia et al., 2024) relies on a tool that parses files into Python-specific class and function structures.
-This fails to perform well in other languages or settings (Yang et al., 2024b) without manual modifications.
-Similarly, the SWE-agent requires modifications to adapt to different tasks (Abramovich et al., 2024; Yang et al., 2024b).
-In contrast, agents designed for PwP are inherently task and language-agnostic due to the expressive action and observation spaces mandated by our environment.
-Moreover, the diverse tasks in PwP-Bench require agents to generalize across a wide range of SWE challenges rather than excel in one narrowly defined area such as resolving pull requests.
-Many existing agents also depend on hand-engineered tools that require human effort to implement and are susceptible to bugs.
-For instance, Agentless (Xia et al., 2024) leverages tools for parsing files into Python-specific structures; CodeAct relies on an IPython kernel (Wang et al., 2024a); SWE-Agent uses dedicated search and file editing tools (Yang et al., 2024a); AutoCodeRover requires a linter (Zhang et al., 2024); SWE-Agent EnIGMA develops specialized tools for CTF-style competitions (Abramovich et al., 2024); and SWE-Bench-MM (Yang et al., 2024b) implements a browser view.
-In PwP, these tools are inherently available within the IDE (as detailed in Table 7), and the agent’s task is to effectively use them rather than being explicitly guided on which tool to use for each specific task.
-Finally, current approaches often blur the line between the agent and the environment, as each agent is designed with its own specified action and observation spaces within a self-created environment.
-Programming with Pixels addresses this issue by unifying existing environments into a single, general-purpose platform on which agents operate.
-This clear separation of environment design from agent design standardizes evaluation and also allows any existing agent to be modeled within our framework, making it an important testbed for both current and future SWE agents.
-2.3 Visual Agents and Computer-Use Agents
-A family of recent multimodal agent benchmarks require agents to operate user interfaces using a predefined, limited set of actions (e.g., new_tab, go_back, click [element id](Koh et al., 2024a; Deng et al., 2023; Zheng et al., 2024) .
-These visual agents typically rely on additional prompting—such as set-of-marks techniques that supply an HTML accessibility tree containing textual and positional information—to overcome their inherent poor visual grounding capabilities (Yang et al., 2023b).
-Despite such aids, these agents often fail when faced with the complex and dense IDE interfaces found in our environment.
-A separate family of computer-use agents (Anthropic, 2024; OpenAI, 2025; Gou et al., 2024) are trained to operate with an expressive action and observation space using primitive operations like clicks and keystrokes, without the need for external accessibility elements.
-However, there is no SWE-specific environment for evaluating and further training these agents.
-PwP fills this gap by providing a unified, expressive IDE platform that challenges computer-use agents with realistic and diverse SWE tasks.
-2.4 Expressive Agent Environments
-Prior work on expressive agent environments has predominantly targeted the web domain (Koh et al., 2024a; Deng et al., 2023), entire operating systems (Xie et al., 2024; Bonatti et al., 2024; Rawles et al., 2023), or other general scenarios (Xu et al., 2024).
-Some of these environments, such as OSWorld (Xie et al., 2024), feature general action and observation spaces similar to ours.
-However, although these benchmarks are capable of expressing a wide range of tasks, they do not focus on the unique challenges inherent to software engineering within an IDE.
-For example, while OSWorld offers a broad set of tasks, it is not specifically designed for SWE, resulting in increased computational overhead.
-Software engineering is a diverse and important domain that merits its own dedicated environment.
-Additionally, we design PwP so that existing tool-based software engineering agents can be readily incorporated into our framework.
-Specifically, we modify the source code of the IDE to open up API calls that let us test current tool-based agents.
-Furthermore, PwP-Bench is tailored specifically for multimodal SWE tasks within an IDE, encompassing activities such as pull-request handling, debugging, and image-based code generation across multiple programming languages.
-We also observe that existing agents built for generic UI control often struggle in the PwP environment, as they must interact with a richer set of tools and achieve precise visual grounding within a complex interface containing a large number of interactive elements.
-We further distinguish PwP from other environments in Table 1.
+Multimodal and Computer-Use Agents.
+Recent works have explored using multimodal LLM agents to operate user interfaces such as web browsers (Koh et al., 2024; Deng et al., 2023; Zheng et al., 2024) and operating systems (Xie et al., 2024; Bonatti et al., 2024).
+Agent designs in these settings fall into two categories: (a) agents with predefined action sets (e.g., new_tab, go_back, click [element id]) that receive auxiliary information such as HTML accessibility trees (Yang et al., 2023a) for visual grounding; (b) pure computer-use agents operating with primitive keyboard and mouse actions, relying solely on screenshots (Anthropic, 2024; OpenAI, 2025; Qin et al., 2025).
+PwP supports evaluating both agent designs.
+Further, existing benchmarks such as OSWorld (Xie et al., 2024), AndroidWorld (Rawles et al., 2025), and WindowsAgentArena (Bonatti et al., 2024) evaluate agents on simple tasks like document editing and calendar management, leaving unclear whether performance on these tasks translates to complex, specialized domains like software engineering.
+PwP-Bench fills this gap by providing the first benchmark specifically designed to test whether computer-use agents can handle software engineering tasks.
+While some prior works explores specialized domains such as game playing (Tan et al., 2024) and a concurrent work explores scientific software (Sun et al., 2025), PwP is the first environment and PwP-Bench the first benchmark systematically evaluating computer-use agents for software engineering, a domain that is particularly noteworthy due to the presence of strong specialized agent baselines.
+Software Engineering Agents.
+Software engineering agents have primarily relied on specialized scaffolding tailored to specific tools, languages, or tasks (Jin et al., 2024; Yang et al., 2024b).
+For instance, Agentless (Xia et al., 2024) uses Python-specific parsers, SWE-agent employs task-specific modifications (Abramovich et al., 2024; Yang et al., 2024b), and others depend on hand-engineered components like IPython kernels (Wang et al., 2024a) or custom browser views (Yang et al., 2024b).
+Our work takes a fundamentally different approach by evaluating whether computer-use agents –which interact through the same visual interface as human developers– can match these specialized agents.
+This also tests whether visual interaction with standard developer tools is sufficient for software engineering or if specialized APIs remain necessary.
+As PwP supports evaluating both designs, it enables direct comparison between computer-use and specialized agents across the diverse tasks in PwP-Bench, establishing a unified platform for understanding the capabilities and limitations of different agent designs.
+We refer readers to Appendix C for a more detailed related work.
 3 Programming with Pixels (PwP)
-Modern software engineering (SWE) requires using multiple programming languages, tools, and modalities.
-Furthermore, it relies on a wealth of tools that required tremendous human effort to create—from linters, to visual code debuggers, to project management tools.
-Motivated by these observations, we create Programming with Pixels, an IDE environment that satisfies two properties: (i) it is expressive, meaning that an agent can perform any task that is achievable through a sequence of primitive operations (e.g., typing or clicking) within an IDE; (ii) an agent has access to any tool implemented within the IDE since using a tool amounts to performing a sequence of primitive actions.
-3.1 PwP environment
-We represent the PwP environment as a partially observable Markov decision process (POMDP). We define the PwP POMDP $\langle S,A,O,T,R\rangle$ as:
-  • $\mathbf{S}$ is the set of states describing the IDE and the operating system context, including open files, active editor panels, and cursor positions.
+Testing computer-use agents (CUAs) on software engineering (SWE) requires an environment that captures the full complexity of modern software engineering, which involves multiple programming languages, tools, and modalities.
+Furthermore, a fair evaluation must provide access to the wealth of tools that human developers use and specialized AI SWE agents have access to, such as linters, visual debuggers, and even project management tools.
+To enable such evaluation, we create Programming with Pixels, an IDE environment that satisfies these two requirements.
+First, it is expressive, meaning that an agent can perform any task that is achievable through a sequence of primitive operations (e.g., typing or clicking) within an IDE, which includes a wide range of software engineering activities.
+Second, an agent has access to any functionality implemented within the IDE, since using IDE functionality amounts to performing a sequence of primitive actions.
+PwP environment.
+We represent the PwP environment as a partially observable Markov decision process (POMDP).
+We define the PwP POMDP $\langle S,A,O,T,R\rangle$ as follows.
+$\mathbf{S}$ is the set of states describing the IDE and the operating system (OS) context, including open files, active editor panels, and cursor positions.
 $\mathbf{A}$ is the action space, encompassing all possible keyboard and mouse events.
@@
 The specific action space varies based on the agent setting, described in (§5).
-  • $\mathbf{O}$ is the observation space. The observation space varies based on the agent setting, described in (§5).
+$\mathbf{O}$ is the observation space.
+The observation space varies based on the agent setting, described in (§5).
 $\mathbf{T}$ is the transition function.
@@
 For instance, after the agent finishes editing code to fix a bug, the environment can run a test suite on the updated files to compute a reward.
-Trajectories in PwP can thus resemble real-world development work: an agent can fix a bug in a repository, use a suggestion tool to help with writing code, or create documentation.
-The IDE and its operating system environment track changes, run tests and return reward signals.
-Table 1: Comparison of PwP with existing environments across different dimensions.
-PwP uniquely combines comprehensive IDE tool support with full multimodal support, general action space, and execution-based evaluation, while maintaining software engineering specificity.
-Multi- General Observation State Tools Execution-Based SWE
-Environment modal Action Space Space Checkpointing Reward Specific
-GAIA (Mialon et al., 2023) Text Limited
-SWE-Bench (Jimenez et al., 2023) Text Limited
-SWE-Bench-MM (Yang et al., 2024b) Text, Image Text Limited
-OpenHands (Wang et al., 2024b) Text, Image Tool Output + Browser SWE
-TheAgentCompany (Xu et al., 2024) Text, Image Tool Output + Browser SWE
-WEBSHOP (Yao et al., 2023) Text Browser
-WEBARENA (Zhou et al., 2024) Text Browser
-VWEBARENA (Koh et al., 2024a) Text, Image Screen Browser
-BrowserGym (Chezelles et al., 2024) Text, Image Tool Output + Browser Browser
-OSWORLD (Xie et al., 2024) Text, Image Screen OS
-WindowsAgentArena (Bonatti et al., 2024) Text, Image Screen OS
-PwP (Ours) Text, Image, Video, Audio Screen All IDE Tools
-3.2 Key Features of Programming with Pixels
+Trajectories in PwP thus resemble real-world development work: an agent can fix a bug in a repository, use a suggestion tool to help with writing code, or create documentation.
+The IDE and OS environment track changes, run tests and return reward signals.
+In addition, we discuss five key features of PwP.
+1.
 Expressive observation and action space.
-A typical approach to building agents is to engineer a set of high-level actions for operations like “open file” or “list symbols in file”, and then engineer an environment that supports each action (Xia et al., 2024; Yang et al., 2024b; Wang et al., 2024b).
-Furthermore, agents receive textual outputs that are manually reformatted for each action.
-The key difficulty is that such engineering does not scale to a large number of actions or to the full range of software engineering tasks, and the agent may be specialized to the observation and action space.
-In contrast, PwP preserves standard screen-based user interaction.
-The agent can navigate IDE menus visually, move the cursor, and press keys.
-This makes the environment expressive: an agent can achieve any task that can be achieved through primitive actions in an IDE.
+PwP provides computer-use agents with an unrestricted environment where they can attempt any software engineering task achievable through an IDE’s visual interface, as humans do.
+Unlike environments with predefined action sets, agents can navigate IDE menus visually, move cursors, and press keys to perform more complex actions.
+2.
 Full Spectrum of Developer Tools.
-A modern IDE offers debuggers, linters, version control, refactoring utilities, integrated terminals, and many extensions.
-In particular, PwP is developed on top of VSCode, which has a rich set of built-in functionalities and extensions.
-Implementing each of these into an agent or environment would require significant human effort.
-However, PwP provides these capabilities out-of-the-box within a single environment.
-Agents can set breakpoints, execute code, use language-specific extensions, review error messages, or run tests in a consistent interface.
+When evaluating computer-use agents on SWE tasks, it is imperative that they have a similar level of access to tools as specialized SWE agents, such as those with custom APIs for debuggers, linters, refactoring utilities, and more (Xia et al., 2024; Yang et al., 2024b).
+PwP provides all these tools through IDE’s visual interface, creating a comprehensive test of whether CUAs can leverage the same rich functionality that specialized agents access through APIs.
+3.
 Multimodality and language agnosticism.
-Because the IDE supports numerous programming languages through extensions and built-in modules, PwP naturally covers tasks across Python, Java, JavaScript, Lean, and more without requiring separate integrations.
-For instance, agents can use the same debugger interface for all languages, or use pre-existing linters provided by IDE extensions.
-Beyond screenshots, the environment provides video streams and audio, though we leave exploring these for future work.
-Rich Feedback and State Access.
-PwP can evaluate performance immediately using testing frameworks or compilation checks.
-For instance, if the agent modifies a file, the IDE can automatically trigger a build, update diagnostics, or run tests, generating real-time feedback.
-In cases requiring deeper inspection—such as verifying that a bug is truly fixed—the environment can reveal file-system changes, process states, or test results.
+CUAs promise generality across tasks and domains.
+Software engineering spans many languages such as Python, Java, JavaScript, Lean, and more, with tasks involving multiple modalities, such as text, images, data files, and PDFs, providing a rigorous test of this generality.
+In PwP, the same CUA must handle code generation, UI development, data science, and theorem proving without task-specific modifications.
+For agents requiring visual grounding support, we modified VSCode’s source code to provide rich DOM trees and Set-of-Marks annotations, ensuring fair evaluation across different CUA architectures.
+4.
+Ease of verification.
+PwP provides direct access to the IDE’s internal state, file system, and OS processes for verification.
+When an agent modifies code, we can run test suites, check compilation, and verify correctness.
+This separation between agent interaction (visual) and evaluation (programmatic) makes it easier to verify task completion and provide other sources of feedback.
+5.
+Future adaptability.
+Computer-use agents are improving rapidly, and so are software engineering agents.
+PwP is designed for future adaptability.
+First, adding new benchmarks is as simple as modifying configuration files.
+Second, PwP’s checkpointing is useful for search and RL training methods.
+Third, PwP’s gymnasium interface (Towers et al., 2024) provides a standard interface for evaluation and development.
+Finally, as agents improve and become capable of using more complex tools, the environment (IDE) would automatically incorporate these without architectural changes.
+This makes PwP an extensible platform for evaluating and developing computer-use agents.
+Infrastructure and Implementation
+PwP is deployed in a secure sandboxed docker environment, running open-source VSCode and a minimal operating system.
+Each container is isolated, preventing interference between experiments, ensuring parallel evaluation and facilitating reproducibility.
+We implement checkpointing for the environment state, which is especially useful for backtracking in search algorithms or training RL agents.
+The environment interfaces to VSCode using four channels for real-time screen capture, DoM information, and customizable configuration such as display, CPU/memory limits, etc.
+However, the complex interaction is abstracted away from the user, as they can simply interact with the environment through gymnasium python API (See Figure 9) and install the environment using a simple pip command.
+We refer to subsection A.3 for more details.
+4 PwP-Bench
+We introduce PwP-Bench, a benchmark containing 15 diverse software engineering tasks that span 14 programming languages and multiple modalities.
+Each task provides agents access to the IDE via the PwP environment.
+The goal of PwP-Bench is to test whether computer-use agents (CUAs) can handle the depth and breadth of software engineering activities.
+Tasks.
+PwP-Bench contains 5400 instances sourced from 13 existing code-generation datasets and 2 newly created by us.
+These tasks are designed to be representative of software engineering activities that take place within an IDE.
+Since the IDE is simply a computer program, in principle, these activities should be achievable by a general-purpose computer-use agent.
+We selected the tasks in PwP-Bench according to three key principles: (1) tasks must require substantial interaction with software engineering tooling, (2) each task should require multiple steps, and (3) the benchmark must cover multiple languages and modalities.
+Accordingly, tasks are grouped into four categories:
+  • Code Generation and Editing: These tasks evaluate the ability to generate and edit code.
+This category includes datasets such as HumanEval for code completion, SWE-Bench (Jimenez et al., 2023) and SWE-Bench-Multilingual (Yang et al., 2025) for resolving pull requests, DSBench for data science tasks (Jing et al., 2024), and Res-Q (LaBash et al., 2024) or CanITEdit (Cassano et al., 2024) for code editing.
+Each dataset benefits from different IDE functionality.
+For example, SWE-Bench can take advantage of debuggers and linters, while DSBench may leverage an IPython kernel and extensions for analyzing large data files.
+Code editing tasks can leverage refactoring utilities and repository searches, covering varied input-output formats and end goals.
+  • Multimodal Code Synthesis: These tasks involve creating code based on input images or other visual data.
+Examples include Design2Code (Si et al., 2024b) for UI development, Chart2Mimic (Shi et al., 2024) for generating Python code from chart images, SWE-Bench-MM (Yang et al., 2024b) for multimodal code editing, and DSBench tasks that rely on images, data files, or PDF documents for data analysis.
+  • Domain-Specific Programming: These tasks focus on specialized fields such as ethical hacking (CTF) (Yang et al., 2023b) and interactive theorem proving (miniCTX) (Hu et al., 2024), which demand significant use and interaction with the IDE’s functionality.
+For example, theorem proving requires continuously inspecting goal states via the IDE, while CTF tasks involve analyzing images, running executables, or installing VSCode extensions (e.g., hexcode readers).
+  • IDE-Specific and General SWE Tasks: Since code generation is only one aspect of software engineering, we introduce two novel task sets that evaluate broader SWE skills.
+The first, IDE Configuration, evaluates an agent’s ability to modify IDE settings such as themes, extensions, and preferences.
+These skills involve substantial interaction with the IDE, and are often a precondition for using IDE functionality such as new extensions.
+The second, which we term General-SWE, targets five different non-code activities: performance profiling, code refactoring, debugging bugs in standard libraries, UI mockup design, and code restoration.
+These tasks target practical software engineering tasks typically absent in conventional benchmarks.
+Full details are in Appendix B.2.
+The distribution of tasks across categories and modalities is shown in Figure 10 in the Appendix.
+Computer-use agents that perform well across these tasks would demonstrate strong potential for automating diverse SWE activities across multiple languages, and working with varied input/output modalities such as text, images, data files, and other data types.
+Furthermore, taking advantage of the functionality provided by the agent’s environment is essential.
+Benchmarking Design and Task Setup.
+All tasks are evaluated within the PwP environment.
+Unlike traditional benchmarks, PwP-Bench presents agents with a realistic IDE environment: each agent receives an initial IDE state $S_{i}$ and an instruction $I$, with the goal to achieve a final state $S_{f}$ evaluated via execution-based criteria (e.g., unit tests).
+Among other capabilities, this setup tests whether CUAs can find relevant information from files, directories, and other resources, which is important for complex software development.
+Furthermore, a task is defined by a simple setup script that defines the initial IDE state, the instructions, and the evaluation logic.
+This makes it easy to add new tasks, allowing PwP-Bench to evolve as new benchmarks or better agents are developed.
+PwP-Bench-Lite.
+Because PwP-Bench contains more than 5400 instances in total, running a full evaluation can be computationally expensive.
+To address this, we also provide PwP-Bench-Lite: a smaller subset of 300 instances.
+This subset preserves the overall difficulty and distribution while ensuring equal representation for each task, thereby making rapid experimentation more accessible.
+5 Evaluating Agents in Programming with Pixels
+We evaluate three distinct agent designs in the PwP environment to understand the capabilities and limitations of computer-use agents for software engineering tasks.
+Computer-use agents.
+Computer-use agents interact with the IDE through primitive actions, i.e., keyboard and mouse inputs, while observing the interface visually through screenshots.
+Each agent operates in a turn-based manner, receiving a screenshot each turn and returning an action to progress toward the goal.
+Since most vision-language models without GUI-specific training struggle with raw pixel coordinates, we incorporate Set-of-Marks (SoM) (Yang et al., 2023a).
+With Set-of-Marks, an agent receives both the raw image and a parsed representation of available interface elements (e.g., buttons, text fields), allowing them to interact via element IDs rather than pixel coordinates.
+This design follows previous works (Xie et al., 2024; Koh et al., 2024).
+Computer-use agents with File/Bash APIs.
+Computer-use agents are augmented with direct access to file-editing and bash commands through text APIs.
+The file-editing APIs include operations such as ‘read file’ and ‘string replace’, while bash operations allow command execution in the terminal.
+Agents receive screenshots only when requested via a screenshot action, rather than automatically each turn.
+This design strictly follows Anthropic’s computer-use implementation (Anthropic, 2024).
+Specialized software engineering agents.
+For comparing how well current computer-use agents perform relative to specialized agents, we evaluate mini-sweagent (SWE-agent, 2024), an agent scaffold specifically designed for software engineering.
+Unlike computer-use agents that interact visually with the IDE, mini-sweagent operates entirely through text APIs.
+For multimodal tasks, it receives required images directly as input in its prompt.
+We chose mini-sweagent due to its near state-of-the-art performance on the widely-used benchmark SWE-Bench, as well as its flexibility for adapting to different programming tasks.
+See Appendix D.1 for implementation details.
+Experimental setup.
+We test multiple models as the parametrization for the two computer-use agent designs.
+Specifically, we test four vision-language models: Gemini-Flash-1.5, Gemini-Pro-1.5, GPT-4o, GPT-4o-mini, and we test five models with UI-specific training: closed-source Claude-3.5 Sonnet, Claude-3.7 Sonnet, Claude-4.0 Sonnet, and open-weights Qwen-2.5-VL and Qwen-GUI-Owl-32B.
+For the mini-sweagent, we test Claude-4.0 Sonnet.
+We keep the experimental setup consistent across all tasks and models: for each task instance, the maximum number of iterations is capped at 20 steps; if the agent either exhausts these steps or issues a stop command, the environment’s final state is evaluated using task-specific metrics (see Appendix B.2 for full details).
+For SWE-Bench related tasks, we further evaluate with a maximum of 250 steps in Appendix F.
+Due to computational and budget constraints, we evaluate on PwP-Bench-Lite, which has 300 task instances.
+5.1 Results and Analysis
+Table 1: Performance Evaluation of Different Agents on PwP-Bench by Task Categories. Best numbers are in bold, and best numbers for computer-use agents are underlined.
+Model Code Generation Multimodal Domain-Specific General Overall
+& Editing Code Generation Code Generation SWE Tasks Avg
+Computer-Use Agents
+Gemini-Flash 0.0% 4.3% 0.0% 0.0% 1.1%
+GPT-4o-mini 0.8% 3.7% 0.0% 2.5% 1.7%
+Qwen2.5-VL-72B 0.0% 4.3% 0.0% 5.0% 1.8%
+GUI-Owl-32B 0.0% 0.0% 0.0% 22.5% 3.0%
+Gemini-Pro 2.5% 5.7% 0.0% 7.5% 3.5%
+GPT-4o 0.8% 12.4% 1.7% 10.0% 5.3%
+Claude-Sonnet-3.5 10.7% 8.3% 5.0% 22.5% 10.5%
+Claude-Sonnet-3.7 11.8% 28.5% 8.3% 27.5% 17.7%
+Claude-Sonnet-4.0 16.0% 38.1% 6.7% 37.5% 22.9%
+Computer-Use Agents with File/Bash APIs
+Gemini-Flash 9.5% 11.7% 8.3% 2.5% 8.9%
+GPT-4o-mini 23.6% 17.6% 15.0% 5.0% 17.8%
+Qwen2.5-VL-72B 13.7% 11.8% 6.7% 7.5% 11.0%
+Gemini-Pro 30.0% 16.7% 3.3% 12.5% 18.8%
+GPT-4o 36.2% 41.9% 28.3% 10.0% 32.6%
+Claude-Sonnet-3.5 47.9% 55.1% 43.3% 22.5% 45.5%
+Claude-Sonnet-3.7 51.9% 58.7% 46.7% 27.5% 49.4%
+Claude-Sonnet-4.0 53.4% 58.6% 43.3% 37.5% 50.7%
+Software Engineering Agents
+MiniSweAgent 49.4% 60.3% 40.0% 37.5% 48.8%
+Figure 2: Example of Successful Use of Live Preview Tool in the UI Replication Task The agent successfully uses the live preview tool in the VSCode browser to compare the UI design it made versus the reference design.
+Figure 3: Example of Successful Use of Tool in the Chart Generation Task The agent can compare the generated chart with the reference chart side by side and refine its code accordingly.
+Table 1 summarizes performance across different agent architectures and base models over the four categories of PwP-Bench (task-wise results are in Table 6).
+As seen in the top half of the table, computer-use agents using only primitive keyboard and mouse actions achieve poor performance, with a maximum overall average of 22.9%.
+This is significantly lower than the software-engineering specific agent mini-sweagent, which achieves 48.8% accuracy.
+We attribute this poor performance primarily to limited visual grounding and an inability to interact effectively with the IDE, particularly for file editing and tool usage; see Section 5.1 for further analysis.
+Among all evaluated models, the Claude computer-use agent performs best, likely because it is specifically trained for UI interactions.
+We found that it can leverage basic IDE tools such as HTML live preview, chart visualization, and file navigation, boosting performance on tasks that require visual understanding and IDE navigation.
+Nonetheless, when the same computer-use agents are granted access to just two text APIs (file editing and bash operations) we observe consistent improvements across all categories, with the maximum average accuracy reaching 50.7%, which, for some tasks is comparable to specialized state-of-the-art agent scores (see Appendix E).
+Interestingly, these operations could theoretically be performed without text APIs, since bash operations could be done through the IDE terminal and file editing through the file editor.
+While we show instances of CUAs attempting these visual operations in Section 5.1, they often make mistakes and are unable to recover from errors.
+However, models still struggle to fully leverage the tooling available in the IDE.
+This is evidenced by poor performance on the ‘General SWE’ category, where tasks often require fewer than ten steps when using appropriate IDE tools.
+We analyze the poor performance on General SWE tasks further in the following sections, confirming that these tasks would become simpler if models could use IDE tooling more effectively.
+Overall, our results show that computer-use agents to have some facility for software engineering, but currently require better visual grounding, tool usage, and planning.
+In the following paragraphs, we analyze these strengths and deficiencies in more detail.
+Claude Computer-Use Agent Demonstrates Basic IDE Tool Proficiency.
+Qualitatively, we found that Claude Computer-Use agent can use basic IDE functionalities, including file explorer navigation, file editing, search, browser-based live preview, and image generation and visualization capabilities.
+Figure 2 demonstrates the agent’s effective use of browser tools in UI replication tasks.
+Similarly, Figure 11 illustrates the agent’s ability to utilize multiple tools while editing specific lines in a repository, relying solely on screenshot observations and primitive keyboard/mouse actions.
+Furthermore, we hypothesize that agents have additional latent abilities to use tools that can be activated through prompting or fine-tuning.
+To investigate this, we examined the project refactoring task (such as symbol renaming) in our ‘General-SWE’ benchmark, where Claude initially achieves 25% accuracy when attempting the task.
+However, when explicitly instructed to use precise tools (such as rename or move to file), its accuracy improves to 75% (see Appendix E).
+Computer-Use Agents Demonstrate Poor Visual Grounding Capabilities.
+Figure 4: Agent Hallucinating Screen Contents The agent correctly mentions, search results are displayed (green), it hallucinates an input field containing “disable import error” (red).
+Figure 5: Agent Misidentifying UI Elements The agent fails to identify the correct input field, typing ‘50’ into the settings search bar instead of the word wrap column setting field (red arrow).
+While, Claude Computer-Use agent is able to use basic IDE tools, we found that in general all current CUAs have significant limitations in visual grounding, i.e., the ability to understand the visual input and take actions on the visual IDE interface.
+We identify three primary failure modes.
+First, the agents can often fail to use the correct UI elements.
+For example, in Figure 5 the agent types in the search bar rather than the settings field, while in Figure 14 the model clicks the wrong location.
+Surprisingly, Set-of-Marks did not resolve these issues; agents would instead select incorrect elements.
+Second, the agents often struggle to comprehend the current UI state, such as linter errors indicated by wavy underlines (Figure 12) or hallucinate screen contents (Figure 4).
+Finally, even when the agent can identify a simple error, such as incorrect indentation, it is often not able to fix the error due to struggling with clicking and typing in the proper locations.
+Furthermore, in Appendix F, models frequently completely ignore the visual state information and instead rely on completely memorized action sequences.
+Quantitatively, we found that 20% and 95% of trajectories have at least one visual grounding error in GPT-4o and Claude Sonnet-4.0, respectively (see Appendix F).
+While grounding has been highlighted as a weak point of computer-use agents in web and OS domains (Koh et al., 2024; Xie et al., 2024), the limitations were primarily observed in models without UI-specific training.
+However, our work shows that even models explicitly trained for UI interaction, such as Claude Computer Use (Anthropic, 2024), exhibit these issues in PwP.
+We hypothesize that the deficiencies come from the IDE being particularly information-dense, as well as potentially not being covered by computer-use training datasets.
+Agents Struggle to Use Advanced IDE Functionality.
+Although the best computer-use agent we tested could use basic IDE functionality, all agents lack the ability to leverage more sophisticated IDE tools.
+Specifically, we can see this through the low performance on the ‘General-SWE’ dataset, which focuses on software engineering activities (e.g., profiling, refactoring, debugging) that can be often completed without direct code edits.
+Although these tasks sometimes require only 4-5 steps when using appropriate IDE tools, agents achieve minimal performance, highlighting substantial room for improvement.
+Furthermore, we observed no successful uses of profilers, debuggers (even when explicitly instructed to) when performing the other tasks in our benchmark (see Appendix E).
+Distribution of Functionality Used by Computer-Use Agents with File/Bash Operations.
+Figure 6: Distribution of file/bash calls vs computer-use interaction for computer-use agents.
+As we observed in Table 1, computer-use agents perform much better when they have access to file and bash API calls, which are based on text inputs and text outputs.
+A natural question is to what extent these are using the visual interface versus relying on text-only APIs.
+We study this in Figure 6, which shows the distribution of file/bash API calls versus computer-use interactions on four representative datasets.
+The figure shows a few interesting patterns.
+First, for HumanEval, agents rely entirely on file APIs.
+This is because HumanEval tasks involve simple function completions that are achievable without IDE interaction.
+The lower performance of pure CUAs on this task (25% compared to 100%) demonstrates their inability to perform basic file editing visually.
+Second, for SWE-Bench-MultiModal, surprisingly there are minimal computer-use interactions, primarily using screenshots to understand the open repository or occasionally attempting to open the built-in browser.
+In contrast, the distribution shifts dramatically for Design2Code, where agents frequently open live preview tools to compare generated designs with reference images, and continuous refining the output (see Figure 2).
+In a similar vein, for VSCode tasks, the agents rely entirely on visual IDE functionality to update settings, install extensions, and edit themes.
+These patterns demonstrate that computer-use agents with file/bash APIs have some ability to choose between visual and API based interactions based on the task requirements.
+On datasets such as HumanEval, their performance improvements stem from bypassing their inability to visually perform edits, instead using text APIs.
+Computer-Use Agents Are Rapidly Improving.
+Figure 7: Performance of Claude-Sonnet Computer-Use Agents over time
+Figure 7 compares the performance of Claude-Sonnet 3.5, 3.7, and 4.0 released over a period of 7 months.
+The line shows steady improvement in pure CUAs, with performance nearly doubling from 10.5% to 22.9%.
+Furthermore, from Table 1 we see that the gap between pure CUAs and CUAs with file/bash operations has steadily decreased from 35.0% to 27.8%.
+These results highlight that while a substantial gap remains, rapid progress is being made and continued improvements may eventually close this gap.
+Leveraging the IDE functionality better would improve performance.
+While a single computer-use agent design can perform non-trivially across a wide variety of tasks, our analysis indicates that these models do not fully exploit domain-specific tools.
+To quantify the potential performance gains if agents could effectively use the IDE, we perform an “assisted” experiment.
+In this experiment, we manually engineered a set of IDE-based tool calls representing commonly used IDE functionalities (e.g., live HTML previews, repository structure, symbol outlines).
+Importantly, each API call is achievable using basic operations in the IDE, meaning that in principle, an agent could learn to perform it.
+See Appendix E for full details.
+Figure 8: Assisted versus Computer-Use Agents
+SWE-Bench Design2Code Chartmimic BIRD (T2 SQL)
+Computer Use Agents 0% 23.5% 2.7% 0%
+CUA + File/Bash 15% 48.1% 25.3% 7%
+Assisted 19% 79.5% 61.6% 17%
+Table 8 summarizes the performance improvements of assisted agents, highlighting an average gain of up to 13.3%.
+These results demonstrate that current CUAs have poor interaction capabilities with complex interfaces, yet there is significant scope for improvement.
+The results also suggest that in the near term, performance gains can be achieved by introducing specialized hand-engineered tools into computer-use agents and incorporating existing agent designs in our PwP environment.
+6 Conclusion
+We introduce Programming with Pixels, an environment designed to evaluate computer-use agents on software engineering tasks.
+We also introduce PwP-Bench, a diverse benchmark of 15 tasks spanning the breadth of software engineering across multiple languages and modalities.
+Our extensive evaluations of nine models reveal that pure computer-use agents relying solely on visual interaction perform poorly, while augmenting these agents with simple file and bash text APIs dramatically improves performance.
+Our analysis pinpoints poor visual grounding and an inability to leverage the rich set of functionality in the PwP environment as primary weaknesses.
+Despite these limitations, our findings show that CUAs are improving rapidly, signaling significant potential.
+PwP establishes software engineering as a natural domain for benchmarking whether generalist computer-use agents can reach specialist-level performance on sophisticated tasks.
+7 Acknowledgements
+We thank Daniel Fried, Graham Neubig, Zora Wang, Saujas Vaduguru, Atharva Naik, Riyaz Ahuja, and Weihua Du for helpful feedback.
+We thank Convergent Research and the OpenAI Researcher Access program.
+This work was supported in part by the National Science Foundation under Grant Nos.
+DMS-2434614 and DMS-2502281.
+Pranjal is supported by SoftBank Group–Arm Fellowship.
+Appendix A Programming with Pixels (PwP) Environment
+A.1 Tools
+Previous methods have proposed use of various hand-engineered tools.
+For a fair comparison, all tools should be accessible in the PwP environment.
+Aas shown in Table 2, PwP natively supports all these tools.
+Table 2: Comparison of Hand-engineered Tools across Methods versus PwP. PwP natively supports all tools.
+Method Hand-engineered Tools Supported in PwP
+Agentless (Xia et al., 2024) File Edit, Repository Structure, File Structure
+CodeAct (Wang et al., 2024a) File Edit, IPython, Bash
+SWE-agent (Yang et al., 2024a) Search File, Search Text, File Edit
+EnIGMA (Abramovich et al., 2024) SWE-agent Tools + Debugger, Terminal, Connection Tool
+swebench-mm (Yang et al., 2024b) SWE-agent Tools + View Webpage, Screenshot, Open Image
+A.2 Comparison with Other Environments
+Table 3: Comparison of different environments across multiple dimensions
+Computer-Use Execution-Based Specialized SWE
+Environment Environment? Reward Domain Specific
+GAIA (Mialon et al., 2023)
+WEBSHOP (Yao et al., 2023)
+WEBARENA (Zhou et al., 2024)
+VWEBARENA (Koh et al., 2024)
+BrowserGym (Chezelles et al., 2024)
+OSWORLD (Xie et al., 2024)
+AndroidWorld (Rawles et al., 2025)
+WindowsAgentArena (Bonatti et al., 2024)
+ScienceBoard* (Sun et al., 2025)
+Cradle* (Tan et al., 2024)
+PwP (Ours)
+In Table 3, we compare PwP with existing environments across multiple dimensions. We evaluate environments along the following dimensions:
+  • Computer-use environment: Whether the environment is designed for computer-use agents, and thereof whether it supports multimodal interaction.
+  • Execution-based evaluation: Use of runtime execution to verify the correctness of agent actions
+  • Specialized: Whether the environment is designed for general and basic tasks, such as web navigation, or is it designed for a more sophisticated, specialized and potetntially economically important tasks.
+Only Cradle (Tan et al., 2024) and ScienceBoard (Sun et al., 2025) are specialized for Game Playing and using Scientific softwares respectively.
+  • SWE-specific: Whether the environment is purposefully designed for software engineering tasks
+Further, ours support other engineering features that others do not.
+For instance, PwP also support streaming video and audio, something other environments do not support out of the box.
+Further, unlike environments such as OS-World, which require manual creation of environment image, PwP is natively docker based, and is based on simple scripts, that can be easily used to modify startup scripts and other configurations for future adaptations.
+Finally, we also specifically suppport state checkpointing which supports storing file system and complete process state, and is especially useful for search-based methods.
 1bench = PwPBench(dataset=’swebench’)
@@
 The code shows basic initialization, action execution, environment control, and reward handling.
-Future Adaptability.
-As agents continue to evolve, PwP provides a unified environment for incorporating new benchmarks and training.
-For example, PwP is amenable to reinforcement learning due to its use of the standard gymnasium (Towers et al., 2024) interface and its checkpointing functionality.
-We show an example of how to interact with PwP in Figure 2.
-Checkpointing also allows for backtracking in search-based methods (Koh et al., 2024b; Putta et al., 2024).
-As software engineering practices progress and new IDE tools emerge, PwP incorporates them without additional engineering overhead.
-Further, as agents become more ubiquitous, it is imperative to evaluate their capabilities in pair-programming scenarios with human developers.
-PwP supports concurrent user-agent interaction, potentially enabling new kinds of pair programming or real-time collaboration studies.
-Finally, adding new tasks requires only minimal modification to configuration and evaluation files.
-3.3 Infrastructure and Implementation
+A.3 Infrastructure and Implementation
 PwP is deployed in a secure sandboxed environment.
@@
 Agents can thus insert code, open new files, or navigate menus using the same actions that a human developer would.
-As shown in 2, a Python API is provided for interaction, following a style similar to common reinforcement learning libraries such as gymnasium (Towers et al., 2024).
+As shown in Figure 9, a Python API is provided for interaction, following a style similar to common reinforcement learning libraries such as gymnasium (Towers et al., 2024).
 The API abstracts away the complexity of container management, benchmark management, and handling observations and actions, allowing researchers to focus on agent development.
 Users can query the environment for the latest screenshot, issue an xdotool command, and receive updated states or rewards.
+Examples of xdotool commands include ‘xdotool mousemove 1000 1200’ and ‘xdotool type ’hello world” and are shown in Figure 9.
 The environment’s container configuration is flexible, allowing for software installations, customizable CPU/memory limits, and display settings (e.g., resolution).
 This versatility is crucial for large-scale evaluation, especially when tasks vary in complexity and resource needs.
-4 PwP-Bench
-We introduce PwP-Bench, a benchmark comprising 15 diverse software engineering tasks that span 8 programming languages and multiple modalities.
-Each task provides agents with access to the complete suite of tools available in the PwP environment.
-The purpose of PwP-Bench is to evaluate how well agents handle a broad range of software engineering (SWE) activities, thereby testing the generality of their code generation and SWE capabilities.
-Tasks.
-PwP-Bench contains 5400 instances covering 15 tasks, sourced from 13 existing code-generation datasets and 2 newly created by us.
-These tasks are designed to be representative of the breadth of software engineering—including tasks beyond conventional code generation to capture real-world complexities—and can be expanded as models excel in newer tasks.
-We followed three guiding principles: (1) tasks should require significant interaction with various SWE tools, (2) each task should necessitate multiple steps to complete, and (3) the overall benchmark should span multiple programming languages and modalities.
-Based on these principles, we collected diverse tasks and grouped them into four categories:
-  • Code Generation and Editing: Evaluates the ability of agents to generate and edit code.
-This category includes datasets such as HumanEval for code completion, SWE-Bench (Jimenez et al., 2023) and SWE-Bench-Java (Zan et al., 2024) for resolving pull requests, DSBench (Jing et al., 2024) for data science tasks, and Res-Q (LaBash et al., 2024) or CanITEdit (Cassano et al., 2024) for code editing.
-Each dataset benefits from different tools.
-For example, SWE-Bench can take advantage of debuggers and linters, while DSBench may leverage an IPython kernel and tools for analyzing large data files.
-Code editing tasks can leverage refactoring utilities and repository searches, covering varied input-output formats and end goals.
-  • Multimodal Code Synthesis: Involves creating code based on input images or other visual data.
-Examples include Design2Code (Si et al., 2024) for UI development, Chart2Mimic (Shi et al., 2024) for generating Python code from chart images, SWE-Bench-MM (Yang et al., 2024b) for multimodal code editing, and DSBench tasks that rely on images or PDF documents during data analysis.
-  • Domain-Specific Programming: Focuses on specialized fields such as ethical hacking (CTF) (Yang et al., 2023a) and interactive theorem proving (miniCTX) (Hu et al., 2024).
-These tasks demand significant interactivity with IDE components.
-For example, theorem proving requires continuously inspecting states via the IDE, while CTF tasks often involve analyzing images, running executables, or installing VSCode extensions (e.g., hexcode readers).
-  • IDE-Specific and General SWE Tasks: Recognizing that code generation is only one aspect of software engineering, we introduce two novel task sets to evaluate broader SWE skills.
-The first, IDE Configuration, evaluates an agent’s ability to modify IDE settings—such as themes, extension installations, and preferences—that are critical for effective tool use in a complex environment.
-The second, which we term General-SWE, targets non-code activities such as profiling, designing UI mockups, managing Kanban boards, and project refactoring.
-These tasks capture essential operational skills typically required by human developers but largely absent from conventional code generation benchmarks.
-Figure 3 shows the distribution of tasks across all categories.
-An agent that succeeds across these tasks demonstrates strong potential for automating a wide range of software engineering activities.
-For instance, PwP-Bench covers Python, Java, JavaScript, HTML, CSS, Bash, SQL, and Lean, and requires agents to work with text, images, data files, and other data types.
-Furthermore, effective interaction with IDE tools is essential.
+Finally, the environment has been tested on three different operating systems: Ubuntu, MacOS, and Windows.
+Appendix B PwP-Bench
 Figure 10: Distribution of tasks in PWP-Bench across four main categories: Code Generation and Editing, Multimodal Code Synthesis, Domain-Specific Programming, and General SWE Tasks.
 The inner ring shows the main categories while the outer ring shows specific datasets and tasks within each category.
-Table 2: Comparison of existing software engineering benchmarks.
-PwP-Bench provides the largest dataset (5400 instances) and uniquely covers all aspects: multiple languages and modalities, real IDE interaction, interactive coding, and both code generation and general software engineering tasks.
-#Instances Multiple Multiple Real IDE Interactive Non-Code Code-Generation
-Benchmark Languages Modalities Env Coding SWE Tasks SWE Tasks
-SWE-Bench (Jimenez et al., 2023) 2K
-SWE-Bench-MM (Yang et al., 2024b) $\leq 1K$
-LiveCodeBench (Jain et al., 2024) $\leq 1K$
-Aider Polyglot (Aider, 2024) $\leq 1K$
-TheAgentCompany (Xu et al., 2024) $\leq 1K$
-VisualWebArena (Koh et al., 2024a) $\leq 1K$
-OSWORLD (Xie et al., 2024) $\leq 1K$
-WindowsAgentArena (Bonatti et al., 2024) $\leq 1K$
-PwP-Bench (Ours) 5.4K
-Benchmarking Design and Task Setup.
-Every task is evaluated within the PwP environment.
-Unlike traditional benchmarks that provide structured, well-formatted context (for example, supplying all relevant schemas in text-to-SQL), PwP-Bench presents agents with an IDE containing a codebase rich in information.
-Specifically, an agent is provided with an initial environment state $S_{i}$ and an instruction $I$.
-The agent’s goal is to update the codebase to satisfy $I$ and transition to a final state $S_{f}$.
-Only $S_{f}$ is evaluated, using execution-based criteria (e.g., running unit tests).
-This setup requires agents to autonomously discover and extract relevant information from files, directories, and other resources—mirroring the challenges faced in real-world software development.
-Many tasks in PwP-Bench require extensive multi-turn interactions and can be time-consuming.
-To support large-scale evaluations, PwP enables parallelized testing in a sandboxed environment, ensuring both security and reproducibility.
-Tasks can also be configured to restrict or partially allow internet access based on experimental needs.
-Furthermore, as software engineering tasks evolve with advancements in model capabilities, our benchmark is designed to grow over time.
-New tasks can be incorporated by creating simple setup scripts that define the IDE’s initial state and evaluation logic, ensuring that PwP-Bench remains modular and adaptable.
-PwP-Bench-Lite.
-Because PwP-Bench contains more than 5400 instances in total, running a full evaluation can be computationally expensive.
-To address this, we also provide PwP-Bench-Lite—a smaller subset of 300 instances, made up of 20 random samples per task.
-This subset preserves the overall difficulty and distribution while ensuring equal representation for each task, thereby making rapid experimentation more accessible.
-5 Agents in Programming with Pixels
-The primary objective of Programming with Pixels is to enable general-purpose SWE agents.
-We evaluate state-of-the-art agents based on vision-language models in our environment.
-Each agent operates in a turn-based manner, receiving a screenshot each turn and returning an action (keyboard or mouse action) to progress toward the goal.
-This design is the same as used in previous works (Xie et al., 2024; Koh et al., 2024a) and we refer them to as computer-use agents.
-In practice, most vision-language models struggle with raw image inputs.
-To mitigate this, we incorporate Set-of-Marks (SoM) (Yang et al., 2023b), in which the agent receives both the raw image and a parse of available interface element (e.g., buttons, text fields).
-The agent then interacts with element IDs instead of raw pixel coordinates.
-We evaluate two categories of computer-use agents. The first category only outputs keyboard and mouse clicks. That is, the agent uses the following observation and action space:
-  • $\mathbf{O}$: a screenshot and set-of-marks annotations.
-  • $\mathbf{A}$: keyboard and mouse clicks with set-of-marks.
-The second category of computer-use agents has access to file and bash commands supplied by the environment through an API.
-Furthermore, a screenshot is received only when the screenshot action is called, instead of receiving it every turn.
-These actions are provided in PwP in a similar design principle as Anthropic computer-use (Anthropic, 2024) which consists of file operations such as ‘read file’, ‘create file’, and ‘string replace’.
-In summary, these agents use the following observation and action space:
-  • O: a screenshot and set-of-marks annotations and text output from tools if used.
-  • A: keyboard, mouse, and file and bash operation actions.
-Finally, in Analysis 6.2, we create a tool-based agent design compatible with PwP.
-Specifically, we provide agents with domain-specific API calls that represent high-level actions (such as getting file structure) instead of UI interactions.
-Each high-level action is implemented as a sequence of low-level actions executed in PwP.
-This setting lets us test current state-of-the-art SWE agent patterns (e.g., Wang et al.
-(2024b)) within the PwP environment.
-6 Experiments
-Table 3: Performance Evaluation of Different Agents on PwP-Bench by Task Categories
-Inputs Outputs Model Code Generation Multimodal Domain-Specific General Overall
-& Editing Code Generation Code Generation SWE Tasks Avg
-Screenshot + SoM Keyboard + Mouse GPT-4o 0.8% 12.4% 1.7% 10.0% 5.3%
-GPT-4o-mini 0.8% 3.7% 0.0% 2.5% 1.7%
-Gemini-Flash 0.0% 4.3% 0.0% 0.0% 1.2%
-Gemini-Pro 2.5% 5.7% 0.0% 7.5% 3.%
-Claude-Sonnet 10.0% 8.2% 5% 20.0% 10.2%
-Screenshot + SoM + Tool Output Keyboard + Mouse + Tool Call (File, Bash) GPT-4o 37.0% 41.9% 28.3% 5.0% 32.3%
-GPT-4o-mini 23.6% 17.5% 15.0% 5.0% 17.8%
-Gemini-Flash 9.5% 11.7% 8.3% 2.5% 8.9%
-Gemini-Pro 30.9% 16.7% 3.3% 5.0% 18.1%
-Claude-Sonnet 52.1% 55.1% 43.3% 20.0% 46.8%
-Figure 4: Example of Successful Use of Live Preview Tool in the UI Replication Task The agent successfully uses the live preview tool in the VSCode browser to compare the UI design it made versus the reference design.
-Figure 5: Example of Successful Use of Tool in the Chart Generation Task The agent can compare the generated chart with the reference chart side by side and refine its code accordingly.
-Experimental setup.
-We evaluate two categories of baseline agents as described in Section 5.
-For each configuration, we test five state-of-the-art vision-language models (VLMs): Gemini-Flash-1.5, Gemini-Pro-1.5, GPT-4o, GPT-4o-mini, and Claude-3.5 Sonnet.
-With the exception of Claude—which is natively trained for UI interaction—the remaining models are provided with SoM.
-At each timestep, an agent receives an observation (with the observation space determined by its category) and returns an action.
-The complete history of observations and actions is incorporated into the model’s context.
-For each task instance, the maximum number of iterations is capped at 20 steps; if the agent either exhausts these steps or issues a stop command, the environment’s final state is evaluated using task-specific metrics (see Appendix B for full details).
-Notably, the agent design remains the same throughout all tasks.
-Due to computational and budget constraints, we evaluate on PwP-Bench-Lite, which has 300 task instances.
-6.1 Results
-Figure 6: Example of the Claude Computer-Use agent successfully using multiple IDE tools to complete a repository level code-editing task.
-Figure 7: Example of Agent Hallucinating Screen Contents While the agent correctly mentions, search results are displayed (green text), it hallucinates an input field containing “disable import error” (red)
-Figure 8: Example of wrong mouse click by Claude-Computer Use Agent The agent attempted to click Settings icon but clicked at the wrong location.
-Table 3 summarizes performance across different agent architectures and base models over the four categories of PwP-Bench.
-As seen in the top half of the table, when using only primitive keyboard and mouse actions most agents have poor performance, with a maximum overall average of 10.2%.
-We attribute this poor performance primarily to limited visual grounding and an inability to interact effectively with the IDE—particularly for file editing and tool usage (see Section 6.2 for further analysis).
-In contrast, when agents are granted access to file editing and bash operations through API calls rather than relying solely on UI interactions, we observe consistent improvements across all categories, with maximum average accuracy reaching 46.8%.
-Among the evaluated models, the Claude computer-use agent performs best, likely because it is specifically trained for UI interactions.
-We found that this agent is able to leverage basic IDE tools—such as HTML live preview, chart visualization, and file navigation—to boost performance on tasks requiring visual understanding and IDE navigation.
-Importantly, we find for the first time that a single computer-use agent can achieve performance comparable to and often surpassing (See Appendix C) state-of-the-art methods across a wide variety of software engineering (SWE) tasks—encompassing multiple languages, modalities, and domains—while operating within a single unified environment and interface, and no specific hand-crafted tools, unlike existing SoTA methods.
-Nonetheless, as detailed in Section 6.2, the models struggle to fully take advantage of the IDE tooling.
-This is evidenced by the poor performance on the ‘General SWE’ dataset, where tasks are often as simple as editing IDE settings and often require fewer than four clicks to complete.
-As we show in Section 6.2, these tasks become simpler if the models could use the IDE tooling more effectively.
-Overall, while the results point toward a promising direction for developing general computer-use SWE agents, significant improvements are still needed in visual grounding, tool usage, and planning.
-We analyze these next.
-6.2 Analysis
-Agents Demonstrate Poor Visual Grounding Capabilities.
-Our qualitative analysis across multiple VLMs on PwP-Bench reveals significant limitations in visual understanding—even for basic IDE interactions.
-We identify two primary failure modes.
-First, models frequently fail to correctly identify the UI elements intended for interaction, as demonstrated in Figure 78.
-In agents using set-of-marks (SoM), this issue manifests as incorrect element selection, while without SoM it leads to inaccurate mouse positioning.
-Second, models struggle to comprehend the current UI state.
-As shown in Figure 910, they consistently fail to recognize highlighted elements, cannot detect linter errors indicated by wavy underlines (Figure 11), and often confuse active panels—resulting, for example, in typing into search bars rather than file editors (Figure 9).
-While similar issues have been documented in web and OS domains (Koh et al., 2024a; Xie et al., 2024), these limitations were primarily observed in models without UI-specific training.
-However, our work, shows even models explicitly trained for UI interaction (Anthropic, 2024), including Claude-Computer Use, exhibit these issues in PwP—likely due to the increased complexity of the IDE interface.
-Agents Fail to Edit Files.
-File editing is a basic capability required in most SWE tasks.
-However, we find that the deficiencies in visual grounding significantly impact the file editing capabilities of current agents that use basic actions (clicking and typing).
-For example, even when provided with cursor location information in textual form, these models struggle to interpret such data amid complex UI elements.
-Models fine-tuned for UI interactions still commit basic editing errors—such as incorrect indentation and text misplacement—and are unable to recover from these errors (see Appendix for examples).
-We speculate these limitations could stem from two factors: (i) model overfitting to user interfaces in their training domains, or (ii) the increased complexity of the PwP IDE interface, which contains substantially more interactable elements than typical web or OS environments.
-Addressing these limitations represents an important direction for future work.
-Although direct file access via tool operations is available, UI-based editing confers unique advantages for tasks such as editing Jupyter notebooks, comparing changes, or modifying specific sections of large files.
-These results underscore two limitations: (i) current VLMs are challenged by complex UI interactions beyond simple web/OS interfaces (Xie et al., 2024; Koh et al., 2024a), and (ii) the inability to effectively perform UI-based editing prevents agents from leveraging valuable IDE features that could have improved their performance.
-Figure 9: Example of Agent Misidentifying UI Elements The agent fails to identify the correct input field, typing ‘50’ into the settings search bar instead of the word wrap column setting field (red arrow).
-Figure 10: Example of Agent Misidentifying Active Panel The agent fails to recognize the active editor panel, incorrectly typing into the search bar (red arrow) instead of the file editor.
-Claude Computer-Use Agent Demonstrates Basic IDE Tool Proficiency.
-Our analysis shows that the Claude Computer-Use agent successfully uses fundamental IDE functionalities, including file explorer navigation, file editing, search, browser-based live preview, and image generation and visualization capabilities.
-Figure 4 demonstrates the agent’s effective use of browser tools in UI replication tasks.
-Similarly, Figure 6 illustrates the agent’s ability to coordinate multiple tools while editing specific lines in a repository, relying solely on screenshot observations and primitive keyboard/mouse actions.
-Second, we hypothesize that the agent has additional abilities to use tools that can be uncovered through prompting or fine-tuning.
-To investigate this, we considered the symbol renaming task in our ‘General-SWE’ benchmark, in which Claude initially achieves 0% accuracy when attempting the task.
-However, when explicitly instructed to use the renaming tool, its accuracy improves to 50% (see Appendix C).
-Agents Struggle to Use IDE Functionality.
-Despite these successes with basic IDE functionality, computer-use agents demonstrate significant limitations when interfacing with more complicated IDE tools.
-We observe no successful instances of debugger usage or symbol listing operations.
-Models without specific UI interaction training struggle even with fundamental tools such as HTML live preview, image visualization, and graph generation capabilities.
-While Claude demonstrates competency with basic tools, it fails to effectively use advanced tools like profilers or debuggers.
-To further evaluate these capabilities, we developed the ‘General-SWE’ dataset, which focuses on software engineering activities (e.g., profiling, refactoring, debugging) that can be completed without direct code modification.
-Although these tasks sometimes require only 4-5 steps when using appropriate IDE tools, agents achieve minimal performance, highlighting substantial room for improvement.
-Training Models to Use IDE Tools Better Would Improve Performance.
-While a single computer-use agent design can perform well across a wide variety of tasks, our results indicate that these models do not fully exploit domain-specific tools.
-As an indication of the potential for performance gains if the agent was able to effectively use the IDE, we perform an “assisted” experiment.
-For the assisted experiment, we manually engineer a set of API calls that are useful for the tasks.
-For example, in Design2Code, the assisted agents have an API call for a live HTML preview, while for SweBench it has API calls for retrieving repository structure and symbol outlines.
-Importantly, each API call is achievable using basic operations in the IDE, meaning that in principle an agent could learn to perform it.
-To ensure that each API call is achievable in the IDE, we implement each API call by executing a fixed sequence of low-level IDE actions, with the details abstracted away from the agent.
-A complete list of tools available in different environments is in Table 7 (see Appendix C).
-Table 4 compares the performance of these assisted agents with that of standard computer-use agents across four datasets for which we manually created tools.
-The assisted agents achieved up to a 13.3% improvement in average scores relative to the non-assisted agents.
-This suggests that training agents to explore and use the built-in IDE functionality would yield performance gains.
-It also suggests that in the near term, we can get performance gains by introducing hand-engineered tools into the computer-use agent and incorporating existing agent designs in our PwP environment.
-Table 4: Comparison of Different Agent Types Across Selected Tasks
-SWE-Bench Design2Code Chartmimic BIRD (T2 SQL)
-Primitive Agents 0% 23.5% 2.7% 0%
-Computer Use 15% 48.1% 25.3% 7%
-Assisted 19% 79.5% 61.6% 17%
-Figure 11: Example of Agent Missing Visual Error Indicators The agent fails to recognize linter error indicators (wavy underlines).
-Figure 12: Example of Agent’s Inability to Perform File Editing The agent incorrectly positions new content in the file editor.
-Agents Are Incapable of Recovering from Errors.
-Next, we find that current agents show limited error recovery capabilities.
-When an action fails to execute correctly, models tend to persistently repeat the same failed action without exploring alternatives.
-Similarly, if an agent selects an incorrect action, it continues along an erroneous solution path without recognizing or correcting the mistake.
-In an experiment designed to probe this behavior, we deliberately suppressed one of the model’s actions.
-Despite the environment’s screenshot clearly showing an unchanged state, the models proceeded with their planned action sequence as though the suppressed action had succeeded.
-This behavior suggests a heavy reliance on memorized action sequences rather than dynamic responses to visual feedback, resulting in exponentially increasing errors and poor performance.
-7 Conclusion
-In this work, we introduce PwP, a unified environment that challenges the prevailing tool-based paradigm by enabling direct interaction with IDEs through basic computer-use actions like typing and clicking.
-This approach allows for a wide range of tasks to be modeled without language- or domain-specific modifications.
-Second, we introduce PwP-Bench, a unification of existing SWE datasets evaluated in PwP.
-We find that general-purpose computer-use agents can approach or outperform previous state-of-the-art results, without any task-specific modifications.
-This suggests that the dominant paradigm of building specialized text-based tools for SWE agents may be superseded by end-to-end computer-use agents.
-However, our analysis reveals that state-of-the-art agents are still incapable of using the extensive set of tools available in PwP, and could perform better if they could use them.
-Our work opens up an exciting new direction of development of computer-use agents for SWE tasks, an important step towards reaching truly general-purpose SWE agents.
-Finally, PwP provides a common platform that supports existing agents and benchmarks, making it a foundation for future research on SWE agents.
-Acknowledgments.
-We thank Daniel Fried, Graham Neubig, and Zora Wang, Saujas Vaduguru, Atharva Naik, Riyaz Ahuja, Weihua Du for helpful feedback, Google and Open AI credit programs, and Convergent Research.
-Appendix A Programming with Pixels (PwP) Environment
-A.1 Tools
-Previous methods have proposed use of various hand-engineered tools. However, as shown in Table 5, PwP natively supports all these tools.
-Table 5: Comparison of Hand-engineered Tools across Methods versus PwP. PwP natively supports all tools.
-Method Hand-engineered Tools Supported in PwP
-Agentless (Xia et al., 2024) File Edit, Repository Structure, File Structure
-CodeAct (Wang et al., 2024a) File Edit, IPython, Bash
-SWE-agent (Yang et al., 2024a) Search File, Search Text, File Edit
-EnIGMA (Abramovich et al., 2024) SWE-agent Tools + Debugger, Terminal, Connection Tool
-swebench-mm (Yang et al., 2024b) SWE-agent Tools + View Webpage, Screenshot, Open Image
-A.2 Comparison with Other Environments
-In Table 1, we compare PwP with existing environments across multiple dimensions.
-We focus on environments designed for software engineering agents as well as those developed for web agents and operating system interaction.
-Since many prior code agent efforts do not explicitly separate their agent and environment architectures, we evaluate their environments based on their published interfaces (e.g., SWE-Bench, SWE-Bench-MM).
-We evaluate environments along the following dimensions:
-  • Multi-modal support: The ability to process diverse input modalities including text, images, video, and audio
-  • General action space: Whether the environment provides a general-purpose interface for tool interaction.
-For instance, OSWorld, WindowsAgentArena, and BrowserGym enable keyboard/mouse interactions through visual inputs, creating a general action and observation space.
-In contrast, environments like SWE-Bench and OpenHands restrict interactions to predefined tool-specific APIs.
-While OpenHands does allow BrowserGym interface, which in turns allows for general action and observation space, the generality is limited to browser interactions and does not apply to other SWE-related tools.
-  • Observation space: The range of environmental states and feedback accessible to an operating agent
-  • State Checkpointing: Native support for preserving file system and process states, a capability unique to PwP
-  • Tools: The set of available tools and capabilities accessible to agents
-  • Execution-based evaluation: Use of runtime execution to verify the correctness of agent actions
-  • SWE-specific: Whether the environment is purposefully designed for software engineering tasks
-While some environments offer general action and observation spaces, they are not specifically designed for software engineering tasks.
-Programming with Pixels uniquely combines SWE-specific design with general-purpose interaction capabilities, enabling the development of general-purpose computer-use agents for software engineering.
-Appendix B PwP-Bench
+Note that the figure is not drawn based on relative size of tasks.
+Table 4: Number of instances for each task in PwP-Bench
+Task Number of Instances Languages
+HumanEval 165 Python
+Design2Code 485 HTML/CSS/JS
+ChartMimic 600 Python
+InterCode 100 Python, Bash
+RES-Q 100 Python
+CanItEdit 105 Python
+VSCode 20 -
+Bird 500 SQL
+DSBench 112 Python
+SWE-bench 2000 Python
+SWE-Bench-Multilingual 91 C++, Typescript, Javascript, Rust, Go, C, Ruby, PhP, Java
+Swebench-$M$ 510 Javascript
+SWT-Bench 276 Python
+Minictx 381 Lean
+General SWE 20 -
+B.1 Tasks
+Figure 10 shows the set of tasks across all categories.
+Further, Table 4 shows the number of instances for each task in the full benchmark, along with the languages used in each of the tasks.
+PwP-Bench-Lite contains 300 instances, which is a random sample of 20 instances from each task.
+B.2 Evaluation
+All tasks are evaluated using programmatic verifiers.
+These verifiers are typically run on an separate environment, not accessible to the agent.
+This typically works based on fetching relevant files and information from the agent environment, and then running through task-specific evaluation scripts on a separate environment.
+However, to the user, this is abstracted away, and they simply have to call ‘env.get_reward()’ to get the exact score or correctness signal based on task.
 Metrics
@@
 While MCQ questions are evaluated using Accuracy, the code generation part is evaluated using linear normalization between the baseline score (of the competition) and the score of the winner of competition.
-Appendix C Results
+VSCode and General SWE Tasks
+In this section, we detail the VSCode and General SWE tasks in PwP-Bench, created by us.
+The VSCode tasks are mostly designed to evaluate the ability of agents to use basic VSCode features, such as renaming all instance of a symbol in file, installing extensions, changing themes, modifying specific settings.
+All these tasks are evaluated based on final IDE state, either by invoking the ‘code’ cli tool, configuration files stored in environment filesystem, or through direct access to VSCode state provided by PwP (see  subsection A.3).
+General-SWE tasks, involves 5 categories of tasks: 1.) QA based on code profiling (evaluated based on final answer by model which requires using appropriate profiling tools), 2.) code refactoring (assessed through automated tests on the final repository state), 3.) debugging bugs in standard libraries (evaluated based on the correctness of final code state), 4.) UI mockup design (assessed using CLIP scores), and 5.) code restoration, where the agent leverages VSCode’s timeline feature to recover corrupted codebases, evaluated by the correctness of the restored state.
+Comparison with Other Benchmarks
+In  Table 5, we further compare PwP-Bench with other existing benchmarks.
+Table 5: Comparison of existing software engineering benchmarks.
+PwP-Bench provides the largest dataset (5400 instances) and uniquely covers all aspects: multiple languages and modalities, real IDE interaction, interactive coding, and both code generation and general software engineering tasks.
+#Instances Multiple Multiple Real IDE Interactive Non-Code Code-Generation
+Benchmark Languages Modalities Env Coding SWE Tasks SWE Tasks
+SWE-Bench (Jimenez et al., 2023) 2K
+SWE-Bench-MM (Yang et al., 2024b) $\leq 1K$
+LiveCodeBench (Jain et al., 2024) $\leq 1K$
+Aider Polyglot (Aider, 2024) $\leq 1K$
+TheAgentCompany (Xu et al., 2024) $\leq 1K$
+VisualWebArena (Koh et al., 2024) $\leq 1K$
+OSWORLD (Xie et al., 2024) $\leq 1K$
+WindowsAgentArena (Bonatti et al., 2024) $\leq 1K$
+PwP-Bench (Ours) 5.4K
+Appendix C Related Work
+C.1 Comparison to Software Engineering Agents
+Task-specific SWE benchmarks
+Early neural code generation approaches were typically evaluated on fixed input-output pairs—for example, generating code from docstrings (Chen et al., 2021) or from general textual descriptions (Austin et al., 2021).
+Subsequent benchmarks extended these evaluations to interactive settings, such as resolving GitHub pull requests or writing unit tests for real-world code repositories (Jimenez et al., 2023; Zan et al., 2024; Mündler et al., 2025).
+More recently, efforts have broadened the scope of code generation to include multimodal tasks, where vision models must interpret images to generate correct code or edits (Si et al., 2024b; Shi et al., 2024; Jing et al., 2024; Yang et al., 2024b).
+However, each of these benchmarks is confined to specific languages, modalities, or task types.
+In contrast, our proposed PwP-Bench unifies these diverse evaluations into a single framework, encompassing multimodal and multilingual challenges that require interaction with a broad suite of IDE tools.
+Using this unified approach we reproduce the performance of established benchmarks and encourage the development of general-purpose agents capable of handling a variety of new software engineering tasks.
+We further compare our work with previous efforts in Tables 3 and 5.
+Software Engineering (SWE) Agents
+Recent work has explored “code agents” that move beyond single-step neural code generation toward interactive methods, where intermediate feedback from tools informs subsequent actions.
+However, many of these approaches specialize in particular tools or programming languages (Jin et al., 2024; Yang et al., 2024b), limiting their broader applicability.
+For example, Agentless (Xia et al., 2024) relies on a tool that parses files into Python-specific class and function structures.
+This fails to perform well in other languages or settings (Yang et al., 2024b) without manual modifications.
+Similarly, the SWE-agent requires modifications to adapt to different tasks (Abramovich et al., 2024; Yang et al., 2024b).
+In contrast, agents designed for PwP are inherently task and language-agnostic due to the expressive action and observation spaces mandated by our environment.
+Moreover, the diverse tasks in PwP-Bench require agents to generalize across a wide range of SWE challenges rather than excel in one narrowly defined area such as resolving pull requests.
+Many existing agents also depend on hand-engineered tools that require human effort to implement and are susceptible to bugs.
+For instance, Agentless (Xia et al., 2024) leverages tools for parsing files into Python-specific structures; CodeAct relies on an IPython kernel (Wang et al., 2024a); SWE-Agent uses dedicated search and file editing tools (Yang et al., 2024a); AutoCodeRover requires a linter (Zhang et al., 2024); SWE-Agent EnIGMA develops specialized tools for CTF-style competitions (Abramovich et al., 2024); and SWE-Bench-MM (Yang et al., 2024b) implements a browser view.
+In PwP, these tools are inherently available within the IDE (as detailed in Table 7), and the agent’s task is to effectively use them rather than being explicitly guided on which tool to use for each specific task.
+Finally, current approaches often blur the line between the agent and the environment, as each agent is designed with its own specified action and observation spaces within a self-created environment.
+Programming with Pixels addresses this issue by unifying existing environments into a single, general-purpose platform on which agents operate.
+This clear separation of environment design from agent design standardizes evaluation and also allows any existing agent to be modeled within our framework, making it an important testbed for both current and future SWE agents.
+C.2 Comparison to General Visual and Computer-Use Agents
+Visual Agents and Computer-Use Agents
+A family of recent multimodal agent benchmarks require agents to operate user interfaces using a predefined, limited set of actions (e.g., new_tab, go_back, click [element id](Koh et al., 2024; Deng et al., 2023; Zheng et al., 2024) .
+These visual agents typically rely on additional prompting—such as set-of-marks techniques that supply an HTML accessibility tree containing textual and positional information—to overcome their inherent poor visual grounding capabilities (Yang et al., 2023a).
+Despite such aids, these agents often fail when faced with the complex and dense IDE interfaces found in our environment.
+A separate family of computer-use agents (Anthropic, 2024; OpenAI, 2025; Gou et al., 2024) are trained to operate with an expressive action and observation space using primitive operations like clicks and keystrokes, without the need for external accessibility elements.
+However, there is no SWE-specific environment for evaluating and further training these agents.
+PwP fills this gap by providing a unified, expressive IDE platform that challenges computer-use agents with realistic and diverse SWE tasks.
+Expressive Agent Environments
+Prior work on expressive agent environments has predominantly targeted the web domain (Koh et al., 2024; Deng et al., 2023), entire operating systems (Xie et al., 2024; Bonatti et al., 2024; Rawles et al., 2023), or other general scenarios (Xu et al., 2024).
+Some of these environments, such as OSWorld (Xie et al., 2024), feature general action and observation spaces similar to ours.
+However, although these benchmarks are capable of expressing a wide range of tasks, they do not focus on the unique challenges inherent to software engineering within an IDE.
+For example, while OSWorld offers a broad set of tasks, it is not specifically designed for SWE, resulting in increased computational overhead.
+Software engineering is a diverse and important domain that merits its own dedicated environment.
+Additionally, we design PwP so that existing tool-based software engineering agents can be readily incorporated into our framework.
+Specifically, we modify the source code of the IDE to open up API calls that let us test current tool-based agents.
+Furthermore, PwP-Bench is tailored specifically for multimodal SWE tasks within an IDE, encompassing activities such as pull-request handling, debugging, and image-based code generation across multiple programming languages.
+We also observe that existing agents built for generic UI control often struggle in the PwP environment, as they must interact with a richer set of tools and achieve precise visual grounding within a complex interface containing a large number of interactive elements.
+We further distinguish PwP from other environments in Table 3.
+Appendix D Experimental Setup and Implementation Details
+D.1 Agent Design
+In addition to the details mentioned in Section 5, we provide more implementation details in this section.
+First, the exact version numbers used for different API models are: gpt-4o-2024-11-20, gpt-4o-mini-2024-07-18, claude-3-5-sonnet-20241022, gemini-1.5-flash-preview-001, gemini-1.5-pro-preview-001, claude-3-7-sonnet-20250219, claude-sonnet-4-20250514.
+For the three Claude models, we use the computer-use variants by passing the ‘computer-use’ beta flag in API calls.
+For open-weights models, we run inference on 8 L40s using vLLM.
+We use temperature=0.3 consistently across models.
+For our main experiments, the number of iterations is set to 20 because: a.) for most tasks, 20 iterations is enough to complete the task, b.) increasing the number to more than 20 would increase the computational cost, and since some models didn’t support caching at the time of running the experiments, the cost grows quadratically, c.) we ran experiments with 250 steps on Claude-Sonnet-4.0 on SWE-bench related datasets (see Appendix F); however, we found no difference in trends.
+D.2 Mini-SweAgent
+For the mini-sweagent, we use Claude-4.0 Sonnet.
+We use the same code as the official source code (SWE-agent, 2024), except that we modify it for multimodal tasks so that the agent receives required images as input in its prompt.
+Appendix E Results
 Table 6 presents comprehensive results for all agent designs across 15 datasets in PwP-Bench.
-For comparison, we include previously reported state-of-the-art results, which represent the best performance achieved on individual datasets through various specialized approaches including finetuned models, custom tool interfaces, specific pipelines, prompts, inference strategies, and verifiers.
-It is important to note that direct comparisons on individual datasets may not provide a complete picture, as state-of-the-art results often utilize different model versions, computational budgets, and domain-specific tools.
-For example, the 55% performance on SWE-Bench was achieved using significantly longer agent trajectories compared to PwP’s 20-iteration limit.
-Given statistical variations and differing evaluation conditions, we emphasize evaluating agents across the full range of tasks rather than focusing on individual dataset performance, as this better aligns with our goal of developing general-purpose SWE agents.
-We make our best effort to include the latest publicly available results, however, there may be minor discrepancies.
-Notably, on the Intercode dataset (Capture the Flag), the computer-use agent significantly outperforms the state-of-the-art method despite the latter (SWE-Agent-Enigma) employing numerous specialized tools for static analysis, dynamic analysis, and networking (See Table 5).
-This improvement is statistically significant (p-value = 0.014, McNemar’s test).
-While the agent did not rely on specific IDE tooling, this result highlights the advantages of a simple, general agent design and suggests the potential benefits of allowing agents to automatically identify optimal tools rather than imposing human priors on the system.
+E.1 Comparison with best reported Specialized SWE Agents
+In this section, we compare computer-use agents with the best reported specialized SWE agents scores on individual datasets.
+In particular, for each dataset, we use 3 different strategies to identify the best reported scores:
+  • Citations: For each dataset, we manually go through the citations and find the most relevant works and look for reported scores.
+  • Official Leaderboard: For some datasets, such as SWE-Bench, we use the official leaderboard to find the best reported scores.
+  • Web-Search Agents: We further prompt ChatGPT-5 thinking to find the latest and highest reported scores on each of the datasets. We then manually verify the results based on the links provided.
+For each dataset, we follow all three strategies and take the highest reported score.
+Typically these results are achieved using specialized approaches including finetuned models, custom tool interfaces, specific pipelines, prompts, inference scaling, and verifiers.
+Therefore, it is important to note that direct comparisons on individual datasets may not provide a complete picture.
+Further, since our evaluations are done on 20 examples from the whole dataset, the results may not be directly comparable.
+Further, while we make our best effort to include the latest publicly available results, there may be still be discrepancies.
+Finally, the search was conducted on 22nd September 2025, and future numbers may change.
+We now list the best reported scores for each dataset:
+  • HumanEval: QualityFlow (Hu et al., 2025) achieves 98.8% performance using Claude-3.5-Sonnet.
+  • SWE-Bench: Highest scores (75.2%) are achieved by a method named TRAE agent (Team et al., 2025), with best reported performance with Claude-4-Sonnet as base model as 74.6%.
+  • SWE-Bench-Multilingual: Highest score publicly reported is 43% (Yang et al., 2025) using Claude-3.7-Sonnet and Swe-agent framework (Yang et al., 2024a).
+  • ResQ: Highest score publicly reported is 58% (LaBash et al., 2024) using Claude-3.5-Sonnet in the official dataset report.
+  • SWT-Bench: Highest score publicly reported is 63.3% (Cassano et al., 2024) using GPT-4o in the official dataset report.
+  • Design2Code: Highest score publicly reported is 90.2% (Si et al., 2024a) using Claude-3.5-Sonnet in the official dataset report.
+  • Chartmimic: Highest score publicly reported is 86.46% using GPT-4o and METAL method (Li et al., 2025). Further they use inference scaling with n=5.
+  • Intercode-CTF: The publicly reported state of the art number is 72% using SWE-Agent-Enigma (Abramovich et al., 2024).
+This is much smaller than the numbers reported by our computer-use agent evaluation, which reaches 100% with the same Claude-3.5-Sonnet model.
+This is surprising, since the method employed numerous specialized tools for static analysis, dynamic analysis, and networking, and we confirmed that the improvement is statistically significant (p-value = 0.014, McNemar’s test).
+  • BIRD: The best reported score is 76.14% (Shkapenyuk et al., 2025) as per the numbers reported in offciail leaderboard.
+  • SWE-Bench-Multimodal: The best reported score is 35.98% using scaffolding over O3, and 34.33% when using OpenHands-Versa (Soni et al., 2025) with Claude-4-Sonnet.
+Overall, the results are often much higher than the numbers achieved by computer-use agents, even with access file and bash APIs.
+Overall, the discussion points out that at present specialized software-engineering agents still perform better, and built scaffolding around computer-use agents might also be helpful.
 Table 6: Performance Evaluation of Different Models Across Task Categories.
-Leged: HE: HumanEval, SB: SWEBench, SJ: Swebench-Java, RQ: ResQ, CI: CaniteEdit, ST: SWTBench, DC: Design2Code, CM: ChartMimic, DS: DSBench, SM: Swebench-MM, IC: Intercode-CTF, BD: Bird SQL, MC: Minictx, VS: VSCode, GS: No-Code SWE Tasks.
-*Much more costly method, ** Estimated Value since only Pass@8 is reported.
+Leged: HE: HumanEval, SB: SWEBench, SJ: Swebench-Multilingual, RQ: ResQ, CI: CaniteEdit, ST: SWTBench, DC: Design2Code, CM: ChartMimic, DS: DSBench, SM: Swebench-$M$, IC: Intercode-CTF, BD: Bird SQL, MC: Minictx, VS: VSCode, GS: General-SWE Tasks.
 Code Generation & Editing Multimodal Domain-Specific No-Code
@@
 Model HE SB SJ RQ CI ST DC CM DS SM IC BD MC VS GS Avg
-Computer-Use Agents   (Screenshot + SoM)
+Computer-Use Agents
+Gemini-Flash 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 15.2% 2.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 1.1%
+GPT-4o-mini 0.0% 0.0% 0.0% 0.0% 5.0% 0.0% 14.8% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 5.0% 0.0% 1.7%
+Qwen2.5-VL-72B 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 17.1% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 10.0% 0.0% 1.8%
+GUI-Owl-32B 0.0% 0.0% 0.0% 0.0% 0.0% 0% 0% 0% 0% 0% 0.0% 0.0% 0.0% 30.0% 15.0% 3.0%
+Gemini-Pro 10.0% 0.0% 0.0% 0.0% 5.0% 0.0% 14.5% 8.1% 0.0% 0.0% 0.0% 0.0% 0.0% 15.0% 0.0% 3.5%
 GPT-4o 5% 0.0% 0.0% 0.0% 0.0% 0.0% 48.7% 0.7% 0.0% 0.0% 5.0% 0.0% 0.0% 20.0% 0.0% 5.3%
-GPT-4o-mini 0.0% 0.0% 0.0% 0.0% 5.0% 0.0% 14.8% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 5.0% 0.0% 1.7%
-Gemini-Flash 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 15.2% 2.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 1.2%
-Gemini-Pro 10.0% 0.0% 0.0% 0.0% 5.0% 0.0% 14.5% 8.1% 0.0% 0.0% 0.0% 0.0% 0.0% 15.0% 0.0% 3.5%
-Claude-Sonnet 20.0% 0.0% 0.0% 15.0% 25.0% 4.2% 18.1% 0.0% 5.0% 10.0% 15.0% 0.0% 0.0% 35.0% 5.0% 10.2%
-Computer-Use Agents   (Screenshot + SoM + File/Bash Operations)
-GPT-4o 85% 25% 15% 30% 50% 17.0% 70.2% 65.5% 11.9% 20% 70% 10% 5% 10% 0.0% 32.3%
+Claude-Sonnet-3.5 20.0% 0.0% 0.0% 15.0% 25.0% 4.2% 18.1% 0.0% 5.0% 10.0% 15.0% 0.0% 0.0% 35.0% 10.0% 10.5%
+Claude-Sonnet-3.7 15.0% 15.0% 0.0% 20.0% 20.0% 0.9% 51.4% 47.6% 0.0% 15.0% 25.0% 0.0% 0.0% 50.0% 5.0% 17.7%
+Claude-Sonnet-4.0 20% 10% 5.0% 20% 20% 20.7% 60.1% 72.4% 10.0% 10.0% 20% 0% 0% 55% 20.0% 22.9%
+Computer-Use Agents with File/Bash APIs
+Gemini-Flash 0.0% 5% 5% 15% 15% 17.1% 19.9% 13.5% 3.2% 10% 25% 0% 0% 5% 0.0% 8.9%
 GPT-4o-mini 60% 10% 5% 20% 30% 16.7% 41.3% 5.5% 8.4% 15% 40% 5% 0% 10.0% 0.0% 17.8%
-Gemini-Flash 0 5% 5% 15% 15% 17.1% 19.9% 13.5% 3.2% 10% 25% 0% 0% 5% 0.0% 8.9%
-Gemini-Pro 85% 10% 15% 15% 40.0% 20.2% 25.6% 24.7% 1.6% 15% 5% 5% 0% 10% 0.0% 18.1%
-Claude-Sonnet 95% 25% 35% 55% 65% 37.4% 83.4% 71.2% 55.7% 10% 100% 15% 15% 35% 5.0% 46.8%
-Previous State of the Art Reported*
-98.8% 55%* 9.9% 58% 63.3% $37.1\%$ 90.2% 71.4% 39.4% 12.2% 72% 30.2% $\approx 25\%$** 35.0% 5.0% 46.8%*
-Appendix D Additional Results
+Qwen2.5-VL-72B 10.0% 5.0% 0.0% 25.0% 25.0% 17.1% 34.1% 13.1% 0.0% 0.0% 5.0% 15.0% 0.0% 15.0% 0.0% 11.0%
+Gemini-Pro 85% 10% 10% 15% 40.0% 20.2% 25.6% 24.7% 1.6% 15% 5% 5% 0% 10% 15.0% 18.8%
+GPT-4o 85% 25% 10% 30% 50% 17.0% 70.2% 65.5% 11.9% 20% 70% 10% 5% 10% 10.0% 32.6%
+Claude-Sonnet-3.5 95% 25% 10% 55% 65% 37.4% 83.4% 71.2% 55.7% 10% 100% 15% 15% 35% 10.0% 45.5%
+Claude-Sonnet-3.7 90% 25% 15% 65% 75% 41.4% 79.2% 81.2% 59.4% 15% 100% 15% 25% 40% 15.0% 49.4%
+Claude-Sonnet-4.0 100% 30% 25.0% 55% 60% 50.6% 86.6% 79.5% 53.1% 15% 100% 15% 15% 50% 25.0% 50.7%
+Software Engineering Agents
+MiniSweAgent 100.0% 25.0% 20.0% 55.0% 65.0% 31.4% 88.1% 80.2% 57.9% 15.0% 90.0% 10.0% 20.0% 55.0% 20.0% 48.8%
+Appendix F Additional Results
+Visual Grounding Errors.
+In Section 5.1, we show that current agents struggle in visual grounding, despite some of these models being specifically trained for visual interfaces.
+To quantify the extent, we manually analyzed 20 random trajectories of two best performing agents: GPT-4o and Claude-3.5-Sonnet.
+In particular, we quantify the number of trajectories where the model had at least one visual grounding error, where a visual grounding error is defined as any of the following: (1) incorrect click, (2) incorrect interpretation of the current state, or (3) interacting with the wrong element.
+Surprisingly, we find that 20% of the trajectories of Claude-Sonnet-4.0, 35% for Claude-3.5-Sonnet, and 95% of the trajectories of GPT-4o contained at least one visual grounding error, indicating significant scope for improving these models for complex visual interfaces such as those demanded by PwP.
 Table 7: Tools available in different environments.
@@
 Training models to use IDE tools better would improve performance.
-In Section 6.2, we demonstrate that models can achieve superior performance when effectively utilizing IDE tools.
+In Section 7, we demonstrate that models can achieve superior performance when effectively utilizing IDE tools.
+In particular, Table 8 shows the performance of assisted agents (averaged across 3 models: GPT-4o, Gemini-1.5-Pro, and Claude-3.5-Sonnet), highlighting an average gain of up to 13.3%.
 However, our analysis reveals two primary limitations in current models’ tool usage: (1) poor visual grounding and inability to handle complex tool interfaces, and (2) failure to prioritize IDE tool-based solutions over manual approaches.
@@
 This operation requires only pressing F2 on a symbol and entering the new name.
-In our evaluation, the Claude agent initially achieved 0% accuracy across four tasks when given no tool guidance.
-However, when explicitly prompted with ”You can utilize the rename feature in VSCode to perform this task,” its accuracy improved to 50%.
+In our evaluation, the Claude agent initially achieved 25% accuracy across four tasks when given no tool guidance.
+However, when explicitly prompted with ”You can utilize the rename feature in VSCode to perform this task,” its accuracy improved to 75%.
 We observed similar patterns across other tasks designed to evaluate tool usage.
@@
 Successful Use of Tools
-We further show a couple of examples of successful tool use in Figure 4 5.
+Figure 11: Example of the Claude Computer-Use agent successfully using multiple IDE tools to complete a repository level code-editing task.
+We further show a couple of examples of successful tool use in Figure 2 3.
 However, we do note that while the agent is able to use the IDE tool through UI interaction, it still may not be able to make optimal use of it as shown in Figure 16.
+Agents Fail to Edit Files.
+File editing is a basic capability required in most SWE tasks.
+However, we find that the deficiencies in visual grounding significantly impact the file editing capabilities of current agents that use basic actions (clicking and typing).
+For example, even when provided with cursor location information in textual form, these models struggle to interpret such data amid complex UI elements.
+Models fine-tuned for UI interactions still commit basic editing errors—such as incorrect indentation and text misplacement—and are unable to recover from these errors (see Appendix for examples).
+We speculate these limitations could stem from two factors: (i) model overfitting to user interfaces in their training domains, or (ii) the increased complexity of the PwP IDE interface, which contains substantially more interactable elements than typical web or OS environments.
+Addressing these limitations represents an important direction for future work.
+Although direct file access via tool operations is available, UI-based editing confers unique advantages for tasks such as editing Jupyter notebooks, comparing changes, or modifying specific sections of large files.
+These results underscore two limitations: (i) current VLMs are challenged by complex UI interactions beyond simple web/OS interfaces (Xie et al., 2024; Koh et al., 2024), and (ii) the inability to effectively perform UI-based editing prevents agents from leveraging valuable IDE features that could have improved their performance.
+Agents Are Incapable of Recovering from Errors.
+Next, we find that current agents show limited error recovery capabilities.
+When an action fails to execute correctly, models tend to persistently repeat the same failed action without exploring alternatives.
+Similarly, if an agent selects an incorrect action, it continues along an erroneous solution path without recognizing or correcting the mistake.
+In an experiment designed to probe this behavior, we deliberately suppressed one of the model’s (Gemini-1.5-Pro) actions.
+Despite the environment’s screenshot clearly showing an unchanged state, the models proceeded with their planned action sequence as though the suppressed action had succeeded.
+This behavior suggests a heavy reliance on memorized action sequences rather than dynamic responses to visual feedback, resulting in exponentially increasing errors and poor performance.
+However, when we repeated the experiment with Claude-Sonnet-4.0, we tested 5 such scenarios, and found only in one case, the agent ignored the screenshot, potentially highlighting that computer-use agents are improving over time.
+Performance on Long Horizon Tasks.
+Table 8: Performance Evaluation of Different Agents on 250 steps on SWE-Bench related tasks.
+Model SWE-Bench SWE-Bench-Multimodal SWE-Bench-Multilingual Average
+Computer-Use Agent 10.0% 30.0% 15.0% 18.3%
+CUA w/ File/Bash Tools 60.0% 30.0% 40.0% 43.3%
+MiniSweAgent 60.0% 30.0% 35.0% 41.7%
+In our main experiments, we had capped the maximum number of agent steps to 20, owing to high cost associated with each of the models.
+However, certain datasets, such as SWE-Bench, typically require much larger number of steps for agent to complete the task.
+In this section, we therefore evluate 3 agents based on Claude-Sonnet 4.0, with 250 steps on 3 relevant datasets: SWE-Bench, SWE-Bench-Multimodal, SWE-Bench-Multilingual.
+The results are shown in Table 8.
+We note, that almost all agents show consistent improvement in performance with higher number os steps.
+However, overall tends remain consistent with 20 steps: Computer-Use Agents with File/Bash APIs show 43.3% performance, and MiniSweAgent shows 41.7% performance, and pure computer-use agents show 18.3% performance.
+Figure 12: Example of Agent Missing Visual Error Indicators The agent fails to recognize linter error indicators (wavy underlines).
+Figure 13: Example of Agent’s Inability to Perform File Editing The agent incorrectly positions new content in the file editor.
+Figure 14: Example of wrong mouse click by Claude-Computer Use Agent The agent attempted to click Settings icon but clicked at the wrong location.
+Figure 15: Example of Agent Misidentifying Active Panel The agent fails to recognize the active editor panel, incorrectly typing into the search bar (red arrow) instead of the file editor.
+Appendix G Qualitative Analysis
+In this section, we consider both positive and negative examples of agent grounding and ability to interact with the complete IDE interface in PwP.
 Figure 16: Performance comparison of GPT-4 agent in Computer-Use and Assisted settings on the ChartMimic dataset.
@@
 The Assisted agent demonstrates superior performance despite seeing the same image but in different context and state.
+Appendix H Discussion
+Computational Overhead of Running PwP
+While PwP provides a much more general interface for software engineering agents, a natural question is what computational overhead it introduces.
+The added computational requirements primarily come from: (1) capturing screenshots using the xdotool library, (2) running the IDE, (3) maintaining a VNC server, and (4) processing video and audio streams via ffmpeg.
+Importantly, only components (1) and (2) are essential for all agents, as video and audio processing are only necessary when agents must interpret visual or auditory cues—a universal requirement for any environment supporting these modalities.
+The VNC server is used solely for debugging or pair programming scenarios and can be disabled when not needed.
+The xdotool commands consume negligible CPU resources (¡¡ 1%) and minimal memory.
+While VSCode does increase memory and CPU utilization, the latency overhead remains limited, and the computational cost is substantially lower than running the large-scale computer-use models that power the agents.
+In summary, despite its comprehensive feature set, the computational overhead of PwP is minimal, with the primary computational demand stemming from the computer-use models themselves rather than the environment.
+Why use IDE over simple Bash Agent?
+While computer-use agents perform worse than even simple API based SWE agents, intutively there still remains a lot of value in utilizing a general interface such as IDE, for software engineering.
+The reason being modern IDEs, have been developed over multiple years of effort, and provide several advantages that are not possible with say bash interface.
+While, theoretically it may still be possible to create equivalent tools, it would take similar tremendous effort, to develop them again for agents, with less reliability.
+To give few examples of myriads advantages of IDEs:
+  • Interactive Debugging Capabilities
+    • IDEs provide rich, stateful debugging interfaces that allow AI agents to set breakpoints, inspect variables, and evaluate expressions dynamically
+    • Unlike CLI debuggers (GDB, LLDB, pdb), IDE debuggers maintain visual context and state, making it easier for AI agents to track program flow and debug complex scenarios
+    • The visual representation of stack traces and variable states is more structured and machine-parseable compared to text-based CLI output
+  • Intelligent Code Refactoring
+    • IDEs maintain a complete Abstract Syntax Tree (AST) of the project, enabling accurate symbol renaming and code restructuring across multiple files
+    • AI agents can leverage IDE’s semantic understanding to perform complex refactoring operations with higher confidence
+    • Unlike text-based search-and-replace in Bash, IDE refactoring tools understand code context and prevent accidental modifications to unrelated symbols
+  • Test Management and Coverage Analysis
+    • IDEs provide structured APIs for test discovery, execution, and result analysis
+    • AI agents can efficiently track test coverage through visual indicators and programmatic interfaces
+    • Real-time test feedback and coverage data is more readily accessible compared to parsing CLI test runner output
+  • Performance Profiling and Analysis
+    • IDE profilers offer structured data about CPU usage, memory allocation, and runtime behavior
+    • Visual representations of performance metrics (flame graphs, memory usage) are easier for AI agents to analyze systematically
+    • Profiling data is available through APIs rather than requiring parsing of complex text-based output
+  • Code Indexing and Semantic Search
+    • IDEs maintain comprehensive code indexes that enable fast, context-aware code search and navigation
+    • AI agents can leverage these indexes for more accurate code understanding and modification
+    • Unlike grep or find, IDE search capabilities understand code structure and can filter based on semantic properties
+  • Extension Integration and Automation
+    • IDE extensions can be programmatically controlled through APIs, allowing AI agents to leverage additional tools seamlessly
+    • Extensions can provide structured data and interfaces that are more reliable for automation compared to parsing CLI tool output
+    • Configuration and coordination of multiple tools can be managed through unified IDE interfaces rather than managing separate CLI tools